### v1

# Measure Twice, Match Once: Flexible Graph Similarity Computation With A Proactive Optimization Strategy

Abstract.
Graph Edit Distance (GED) is an important similarity measure in graph retrieval, which quantifies the minimum cost of transforming one graph into another through edit operations, and offers flexibility by allowing customizable operation costs. Recent learning-based approaches approximate GEDs with the distances between representations in vector spaces. However, these methods often struggle with varying operation costs due to neglecting the impact of these costs on determining optimal graph mappings. Furthermore, they rely on isolated node distances as guidance, necessitating inefficient reactive refinements of mappings. To address these issues, we propose Graph Edit Network (Gen), a novel learning-based approach for flexible GED computation. By identifying the limitations of existing methods in capturing flexibility of GED, we introduce a principled yet simple solution that incorporates the operation costs before establishing mappings. To improve matching efficiency, we propose a strategy that proactively optimizes guidance from a graph perspective. This strategy initializes guidance as each node’s alignment difficulty and captures the interdependencies between matches within and across graphs through a difficulty propagation mechanism, enabling more informed decisions. As a result, Gen selects optimal matches in a single step, minimizing the need for costly refinements. Results on real-world and synthetic datasets demonstrate the effectiveness, time efficiency, and adaptability of Gen, achieving up to 37.8% error reduction and 72.7% inference time reduction compared with state-of-the-art models, while performing robustly under varying cost settings and graph sizes.
graph similarity computation, graph edit distance, neural networks
ccs: Information systems Similarity measuresccs: Computing methodologies Neural networks
Figure 1. The black dashed lines are matches, and the red markings are edits. $f_{1}$ and $f_{2}$ are mappings. $f_{1}$ incurs 2 node substitutions (ns), 1 node insertion, 1 edge deletion, and 2 edge insertions (ei), denoted as $[2,1,0,1,2]$, while $f_{2}$ incurs $[4,1,0,0,1]$. When ns costs 2, other operations cost 1, $f_{1}$ is optimal with a cost of 8, while $f_{2}$ costs 10. However, if ei costs 2, others cost 1, $f_{1}$ costs 8 while $f_{2}$ costs 7, making $f_{2}$ the optimal choice.
1. Introduction
Graph similarity computation is crucial for graph retrieval systems (Zheng et al., 2014; He and Singh, 2006; Wang et al., 2010; Zhu et al., 2012; Zheng et al., 2013; Chang et al., 2023), as it enables the discovery of relevant patterns and relationships. One key measure of graph similarity is Graph Edit Distance (GED), which quantifies the minimum cost of transforming one graph into another through operations, such as node/edge insertion, deletion, or substitution. The flexibility of GED lies in its ability to assign customizable costs to each operation, enabling similarity measures tailored to specific applications. For example, in graph database for recommender systems, higher costs can be assigned to nodes/edges representing key user preferences. This prioritizes the retrieval of graphs that preserve these preferences, which can be used to enhance personalized recommendations.
Conventionally, GED is computed using A*-like algorithms (Abu-Aisheh et al., 2015; Zeng et al., 2009; Chang et al., 2023), which sequentially establish node/edge mappings between graphs. The operation costs assigned to individual operations are crucial in this process, as they directly influence matching decisions. This leads to varying optimal mappings, ultimately affecting the final GED, as illustrated in Fig. 1. However, such a process requires exploring all possible mappings to find the optimal ones with the minimum total cost, resulting in significant latency. Recent studies show that state-of-the-art algorithms struggle to compute exact GED for small graphs with 16 nodes within a reasonable time (Blumenthal and Gamper, 2020). This limits their practicality for modern graph retrieval tasks that require the timely processing of large volumes of graphs.
Recently, with the great success of Graph Neural Networks (GNNs) (Kipf and Welling, 2017; Xu et al., 2019; Veličković et al., 2018; Brody et al., 2022), numerous GNN-based approaches have been proposed (Li et al., 2019; Qin et al., 2021; Zhuo and Tan, 2022; Ranjan et al., 2022; Bai et al., 2019, 2020; Zhang et al., 2021; Doan et al., 2021; Bai and Zhao, 2021; Liu et al., 2022; Piao et al., 2023; Jain et al., 2024; Jia et al., 2023). By leveraging GNNs to generate expressive node representations that capture rich neighborhood information, they effectively approximate GED in the embedding space. These methods are promising alternatives for graph retrieval tasks, where response speed is prioritized over exact mappings. Depending on how node representations are utilized, they can be divided into two categories: graph comparison methods and node comparison methods. Graph comparison methods (Li et al., 2019; Qin et al., 2021; Zhuo and Tan, 2022; Ranjan et al., 2022; Bai and Zhao, 2021) aggregate node representations into a single graph embedding and compute GED by measuring the distance between these embeddings. While computationally efficient, this aggregation leads to a loss of node-level details, which are crucial for considering edit operations. Consequently, these approaches may fail to capture the flexibility inherent in GED computation. In contrast, node comparison methods (Bai et al., 2019; Li et al., 2019; Bai et al., 2020; Zhang et al., 2021; Doan et al., 2021; Piao et al., 2023; Jain et al., 2024; Liu et al., 2022) compare node representations between graphs to compute node matching costs. These costs guide the iterative refinements of mappings by down-weighting high-cost matches for each node. The total cost of the final mappings is then used to approximate the GED. By focusing on node-level differences, similar to conventional algorithms, methods in this category may offer better flexibility.
However, these node comparison methods face two major limitations: (1) Neglect the Role of Operation Costs in Mapping Determination. Most methods fail to account for operation costs. GraphEdX (Jain et al., 2024) attempts to address this by initially establishing mappings without considering operation costs, and then using these costs to reweigh the total costs of mappings afterward. However, operation costs directly influence matching decisions, meaning mappings should vary according to these costs, not just the final GED value. This oversight leads to static mappings under varying cost settings, resulting in inaccurate GED predictions. (2) Reliance on Insufficient Node-level Guidance. These methods use node matching costs as guidance to establish mappings. However, these costs struggle to effectively guide the matching process because they capture only isolated node-level similarities. GED computation, in contrast, requires comparing the interdependencies of matched nodes within graphs to quantify and minimize structural mismatches between graphs after alignment. Node-level guidance lacks this necessary graph perspective. As a result, existing methods rely on a reactive mapping refinement strategy, which improves mappings over multiple rounds with additional constraints, such as considering local node connections (Li et al., 2019; Bai et al., 2020) or minimizing graph transport costs (Doan et al., 2021; Piao et al., 2023; Jain et al., 2024). These refinements significantly increase computational overhead, leading to an inefficient matching process.
To address these limitations, we propose Graph Edit Network (Gen), a novel learning-based model for flexible GED approximation. Gen introduces two key features: (1) Flexible Adaptation to Cost Settings. We identify a key limitation of existing methods in capturing the flexibility of GED: they either neglect or treat operation costs as irrelevant to mapping determination. To address this, Gen incorporates operation costs prior to establishing mappings. This solution aligns with the principle in GED computation, where mappings are dynamically determined based on operation cost, enabling flexible mapping determination across varying cost settings. (2) Effective Graph Perspective Guidance. Unlike existing methods that rely on insufficient node-level guidance and require reactive multi-round refinements to improve mappings, we propose a strategy that proactively optimizes guidance from a graph perspective, thereby minimizing the need for such refinements. First, this strategy initializes guidance as each node’s alignment difficulty based on the global cost ratios of its associated matches. Such guidance provides graph-level comparability and mitigating the risk of myopic, locally optimal decisions. Second, it uses a Difficulty Propagation mechanism to propagate these difficulties within and between graphs, capturing the interdependencies of nodes involved in matches. Within graphs, it respects the node connections, and estimates the alignment difficulty for each neighborhood, broadens the decision-making scope from individual nodes to their neighborhoods. Between graphs, it considers the neighborhood alignment difficulties of involved nodes for each match, facilitating more balanced match selections. Leveraging this guidance, Gen selects optimal matches in a single step, ensuring an efficient matching process. The main contributions are summarized as follows:
  • We propose Graph Edit Network (Gen), which enables flexible GED computation under varying cost settings and employs an effective proactive guidance optimization strategy.
  • We introduce a principled yet simple solution to model the inherent flexibility of GED computation, addressing a gap in graph similarity learning.
  • We propose a proactive optimization strategy with a difficulty propagation mechanism that provides effective graph perspective guidance, allowing optimal match selections in a single step and ensuring an efficient matching process.
Empirical results on three commonly used GED datasets validate the efficiency of the proposed method, achieving up to 37.8% error reduction and 72.7% inference time reduction compared with state-of-the-art models. Gen demonstrates robustness across varying cost settings, consistently delivering strong performance. The ablation study highlights the importance of our proposed strategy, and the hyperparameter sensitivity analysis provides insights into Gen’s behavior. Additionally, experiments on larger synthetic graphs highlight Gen’s reliable performance across different node sizes, confirming its practical utility in real-world applications.
2. Related Work
Graph similarity computation (GSC) research approximates domain-agnostic measures, such as graph edit distance (GED) (Bai et al., 2019; Li et al., 2019) or maximum common subgraph (MCS) (Roy et al., 2022), using neural networks. These GSC methods can be categorized into end-to-end approaches that rely on GNNs and hybrid models that combine GNNs with A* algorithms (Wang et al., 2021; Yang and Zou, 2021; Liu et al., 2023). This paper focuses on end-to-end GED approximation, where GSC methods fall into two main categories: graph comparison methods and node comparison methods.
Graph comparison methods rely on graph-level similarity, with representative methods including (Li et al., 2019; Qin et al., 2021; Zhuo and Tan, 2022; Ranjan et al., 2022; Bai and Zhao, 2021). GMN-emb (Li et al., 2019) generates graph embeddings using a gated mechanism to capture node importance. To improve efficiency, EGSC (Qin et al., 2021) employs a teacher network to fuse graph pair embeddings for contextual predictions, while using a student network to distill them into separate embeddings for storage. Similarly, Eric (Zhuo and Tan, 2022) incorporates a node-graph alignment constraint during training and decouples this module during inference to reduce response time. To model the triangle inequality of the metric, Greed (Ranjan et al., 2022) uses the lower and upper bounds of GED computed by inexact algorithms to frame the distances between graph representations. To improve the interpretability, TaGSim (Bai and Zhao, 2021) predicts graph edit vectors (GEVs) to represent the counts of each operation type by comparing graph representations at multiple scales. Among these methods, while TaGSim might appear capable of capturing flexible GED computation, such a graph comparison method is insensitive to mapping variations. Reweighing its GEVs based on operation costs, as in GraphEdX (Jain et al., 2024), can only lead to inaccurate GED predictions.
Node comparison methods focus on node-level similarities, with representative works including (Li et al., 2019; Bai et al., 2019, 2020; Doan et al., 2021; Piao et al., 2023; Jain et al., 2024). To enhance graph similarity scores, SimGNN (Bai et al., 2019) introduces histogram features based on node similarities. However, these non-differentiable features limit the method’s learning capability, as they cannot participate in training. Except for SimGNN, other approaches adopt a reactive mapping refinement strategy. GMN-match (Li et al., 2019) repeats a fusion-propagation process, starting by fusing representations of cross-graph node pairs using matching-cost-based local attention. The fused representations are then propagated within graphs and adjusted according to local node connections. However, node-level fusion ratios can lead to myopic, locally optimal decisions that directly affect subsequent updates, increasing the risk of error propagation. GraphSim (Bai et al., 2020) reorders multi-scale similarity matrices using BFS traversal to consider local node connections. However, this traversal oversimplifies the complex graph structures, potentially leading to inconsistent performance. Recent methods adopt combinatorial optimization techniques, using the transport cost of established node mappings to approximate GED. GOTSim (Doan et al., 2021) augments the similarity matrix to account for node deletion and insertion operations, then uses a linear assignment solver to align nodes at multiple scales. GEDGNN (Piao et al., 2023) and GraphEdX (Jain et al., 2024) use Gumbel-Sinkhorn networks to iteratively refine mappings. GEDGNN computes both a similarity matrix and a matching matrix to enhance performance. GraphEdX (Jain et al., 2024) designs distance measures for each operation type to establish mappings and subsequently reweighs transport costs with operation costs. However, these optimization-based methods treat graph combinatorial optimization in a matrix formulation, overlooking the node connections. As a result, the cost of their established mappings may fail to reflect the true structural differences between graphs. Furthermore, all methods in this category overlook the role of operation costs in determining mappings, leading to static mappings under varying cost settings.
3. Background
Figure 2. Overview of our proposed Graph Edit Networks (Gen). The dashed node represents the dummy node.
[…]
3.1. Problem Setup
Given a source graph $S$ and a target graph $T$, where the graph with fewer nodes is padded with dummy nodes to ensure that $|\mathcal{V}_{S}|=|\mathcal{V}_{T}|$. The GED computation can be viewed as an error-correcting graph matching problem (Bunke, 1997), which constructs a bijective node mapping $f:\mathcal{V}_{S}\mapsto\mathcal{V}_{T}$ that minimizes the total matching cost. The solution space of such mappings can be represented as a complete bipartite graph $\mathcal{B}=(\mathcal{V}_{S}\cup\mathcal{V}_{T},\mathcal{E}_{\mathcal{B}})$, where each edge $m\in\mathcal{E}_{\mathcal{B}}$ signifies a candidate source-target node match.
Matching Cost. Elements across the graphs are matched by applying edit operations to correct the inherent differences, such as label mismatches. Such corrections incur matching costs $\text{c}(i_{s},i_{t})$, representing the operation cost-weighted differences, as follows.
(1) $\text{c}(i_{s},i_{t})=\text{cost}(op_{(i_{s},i_{t})})\cdot d(i_{s},i_{t})$
Where $\text{cost}(\cdot)$ is a non-negative, customizable cost setting function that assigns an operation cost to each edit $op_{(i_{s},i_{t})}$. $d(i_{s},i_{t})$ quantifies the attribute and structural differences between elements $i_{s}\in S$ and $i_{t}\in T$, which is typically 0 or 1 for nodes/edges, as they are individual elements. For simplicity, $d(\cdot,\cdot)$ is omitted in such cases.
Graph Edit Distance (GED). Given any node mapping $f$, the corresponding operation sequence can be easily deduced, and vice versa. The total cost of $f$ and GED can be computed as follows.
(2) $c(S,T)_{f}=\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\text{c}(s,t)+\sum_{e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}}\text{c}(e,\epsilon)+\sum_{e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})}\text{c}(\epsilon,e^{\prime})$
(3) $\text{GED}(S,T)=min(c(S,T)_{f}|f\subset\mathcal{E}_{\mathcal{B}})$
Where $\epsilon$ represents nonexistent edges, $\text{cost}(s,t)$ signifies the matching cost between $s$ and $t$, $e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}$ refers to an edge $e(s,s^{\prime})$ in $S$ without a correspondence in $T$, and similarly for $e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})$.
Limitation for Flexible GED Computation. When $\text{cost}(\cdot)$ in Eq. 1 assigns a uniform cost (e.g., 1), to each operation, the total cost corresponds to the number of applied operations. Suppose $f^{*}$ is an optimal mapping under this setting. When $\text{cost}(\cdot)$ varies, $c(S,T)_{f^{*}}$ changes accordingly, causing previously optimal mappings to become suboptimal, as illustrated in Fig. 1. Existing methods either neglect or treat operation costs as irrelevant when determining the mappings, resulting in static mappings based solely on inherent element differences and failing to capture the flexibility of GED. Consequently, these methods are incapable of performing customizable searches that adapt to varying cost settings. Furthermore, their reliance on static mappings constrains both theoretical and practical understanding of GED computation.
Problem Formulation. We use the total matching cost $c(S,T)$ of the soft node mapping to approximate GED values, which ranges from $[0,+\infty)$. The total mapping cost is distributed across selected matches. Specifically, each candidate match $m$ in $\mathcal{B}$ is associated with a matching cost, reflecting the cumulative cost needed to transform $\mathcal{N}(s)$ into $\mathcal{N}(t)$. A compensation approximator $g(\cdot)$ is introduced to adjust for duplicated edge cost calculations caused by overlapping neighborhoods. Furthermore, each match is assigned a certainty $\mathbf{1}_{s,t}$ to indicate its soft selection or rejection. The total mapping cost is then expressed as follows.
(4) $c(s,t)=g\bigg{(}\sum_{op\in|\text{op}|}cost(op)\cdot d(\mathcal{N}(s),\mathcal{N}(t))_{op}\bigg{)}$
(5) $c(S,T)=\sum_{s\in|\mathcal{V}_{S}|,t\in|\mathcal{V}_{T}|}\mathbf{1}_{s,t}\cdot c(s,t)$
Where $c(s,t)$ is the matching cost, $cost(op)$ is the operation cost for $op$ applied between $s$ and $t$. For each type $\in\text{op}$ op, $d(\cdot,\cdot)_{op}$ quantifies the differences between $\mathcal{N}(s)$ and $\mathcal{N}(t)$ corrected by $op$.
4. The Proposed GEN Architecture
The overview of our proposed Gen is illustrated in Figure 2. Given a source-target graph pair and customized cost per operation type for each possible match as input, Gen begins by generating node representations for both graphs using a shared GNN Encoder. To enable flexible GED computation, Gen dynamically incorporates operation costs with node distances within the embedding space during the Node Matching Cost Estimation step. Next, Gen proactively optimizes guidance from a graph perspective to minimize the need for reactive refinements. It begins by initializing the guidance as each node’s alignment difficulty to enable globally comparability. Specifically, it first normalizes the matching costs across all potential node matches as global cost ratios, then computes each node’s alignment difficulty as the total cost ratios of its associated matches. To capture the interdependencies of matches, it propagates these difficulties within and between graphs through the Difficulty Propagation mechanism. Within each graph, it respects node connections and estimates the alignment difficulty of each node’s neighborhood. Between graphs, Gen calculates the certainty for each match by considering the neighborhood alignment difficulties of its involved nodes, thus balancing match selection. Finally, the overall matching cost of a graph is determined by summing the certainty-weighted matching costs.
4.1. GNN Encoder
Given a graph as input, Gen uses a $k$-layer GNNs to generate representations for each node based on its $k$-hop local structure. The node representation for node $v$ is initialized as the one-hot encoding of the node label, $\mathbf{x}_{v}\in\mathbb{R}^{|\mathcal{X}|}$, which $|\mathcal{X}|$ is the number of label types. Nodes in the unlabeled graphs can be regarded as having an identical label. We pad a smaller graph between $S$ and $T$ with dummy nodes to ensure $|\mathcal{V}_{S}|=|\mathcal{V}_{T}|$. The representations of dummy nodes are initialized as zeros. We first smooth $\mathbf{x}_{v}$ with a Linear layer to obtain $\mathbf{x}_{v}^{0}\in\mathbb{R}^{d_{hidden}}$. Based on the adjacency matrix $\mathcal{A}$ of graphs, GNNs update node information for each node with aggregated information from the neighborhoods. The encoding process at layer $l$ for node $v$ can be represented as follows.
(6) $\mathbf{x}^{l}_{v}=\psi\bigg{(}\mathbf{x}^{l-1}_{v},\phi(\mathbf{x}^{l-1}_{u}:\mathcal{A}(v,u)=1)\bigg{)}$
Where $\phi$ is an aggregation function, and $\psi$ is a combination function, typically permutation-invariant operations like Sum, Mean, or Max. $\mathbf{x}^{l}_{v}$ and $\mathbf{x}^{l-1}_{v}\in\mathbb{R}^{d_{hidden}}$ represent the output of node $v$ from layers $l$ and $l-1$, respectively. $\mathbf{x}^{l-1}_{u}\in\mathbb{R}^{d_{hidden}}$ represents the output of $v$’s neighbor at the previous layer. The final layer representation $\mathbf{x}^{k}_{v}$ is then passed through a Multilayer Perceptron (MLP) to reduces the dimensionality of the representation to $d_{out}$. By altering GNNs with the ones considering edge information, this method can be extended to edge-labeled graphs without losing generalization.
4.2. Node Matching Cost Estimation
[…]
(7) $\begin{gathered}\text{dist}(s,t)=\mathbf{h}_{s}-\mathbf{h}_{t}\\ c(s,t)=\text{Estimator}\bigg{(}\text{Concat}(cost_{(s,t)},\text{dist}(s,t))\bigg{)}\end{gathered}$
The input cost $cost_{(s,t)}\in\mathbb{R}^{|op|}$ represents predefined costs per operation type between $s$ and $t$. The concatenation of $cost_{(\cdot,\cdot)}$ and $\text{dist}(\cdot,\cdot)$ is fed into the $\text{Estimator}(\cdot)$, an MLP, which consists of fully connected layers, meaning all input entries are connected to each output. This allows the operation costs and the corrections made by the corresponding operations to be combined in the outputs through learnable weights and biases. Additionally, this $\text{Estimator}(\cdot)$ also serves as a compensation function $g(\cdot)$ to adjust the node matching costs for duplicated edge costs by decoding the neighborhood interactions captured within the node embeddings. While more sophisticated compensation methods could potentially enhance accuracy, this has neither been discussed nor explored in the learning-based GED computation literature. In this work, we prioritize efficiency and consider it an open area for future research. Gen is the first work to consider operation costs before the matching process, correcting the misinterpretation of the role of operation costs in previous methods. This simple solution ensures that, while inherent differences between compared elements remain consistent, they can be dynamically adjusted to accommodate varying cost settings. It is theoretically in line with the principle of GED computation (Eq. 1), enhancing the model’s understanding of GED computation. Experiment 5.2.3 further highlights the limitations of existing methods on this issue and validates the effectiveness of our solution, enabling flexible graph similarity computation.
4.3. Proactive Optimization Strategy
Unlike previous work that relies on the insufficient node-level guidance of matching costs to reactively refine $\mathbf{1}_{s,t}$ in Eq. 5. Gen proposes a proactive optimization strategy that offers guidance from a graph perspective to minimize the need for refinements. Gen first computes the global cost ratio for each match, then uses the total cost ratios of each node’s associated matches to compute the node’s alignment difficulty, which serves as initial guidance. Next, Gen incorporates the interdependencies of matches within and between graphs through a difficulty propagation mechanism.
4.3.1. Guidance Initialization
Global Cost Ratio Computation.
Matching costs can vary significantly across matches. While the absolute cost of each match may be clear, comparing these costs directly can be misleading, as some matches might seem optimal for one node but not contribute optimally to the overall graph alignment, leading to myopic, locally optimal decisions for individual nodes. To address this, Gen proposes first normalizes matching costs across all potential node matches as global cost ratios, enabling global comparability:
(8) $\gamma(s,t)=\frac{exp(c(s,t))}{\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}exp(c(s,t))}$
Where $\gamma(s,t)\in\mathbb{R}^{d_{out}}$ represents the cost ratio of a match compared with the overall cost of all candidate matches. Intuitively, a match with a higher matching cost can have a larger ratio, reflecting its global significance within the graph-level alignment.
Node Alignment Difficulty Computation.
Each node is involved in different candidate matches, forming a spectrum of differences between the node and the other graph. Gen gathers the cost ratios for each involved source and target node to compute the node alignment difficulty in the global context as follows.
(9) $\tilde{\gamma}(s)=\sum_{i\in\mathcal{V}_{T}}\gamma(s,i)\quad\text{and}\quad\tilde{\gamma}(t)=\sum_{i\in\mathcal{V}_{S}}\gamma(i,t)$
Intuitively, the higher $\tilde{\gamma}(\cdot)$ means the node is involved in high-cost matches, suggesting that accurate matching of this node is crucial, or it may contribute more to the overall matching cost. This guidance provides a comprehensive assessment of each node’s relative importance across the graphs, helping prioritize nodes that are more critical for global alignment.
4.3.2. Difficulty Propagation
Intra-graph Consideration
Matches within a graph are interdependent due to node connections. Matching one node can constrain the possible matches for other nodes, as the neighborhood structure must be preserved to minimize differences. Similarly, the alignment difficulties of neighboring nodes are important. A node with high alignment difficulty can impose constraints on its neighbors, necessitating careful handling of their alignment. To capture such intra-graph relationships, Gen respects the node connections within the graphs and considers the alignment difficulty of each node’s neighborhood. This broadens the decision-making scope from nodes to neighborhoods. Specifically, Gen propagates the alignment difficulty of each node within its own graph as follows.
(10) $\hat{\gamma}(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Sum}(\tilde{\gamma}(v),\text{Sum}(\tilde{\gamma}(u):\mathcal{A}(v,u)=1))\bigg{)}$
Here, $\hat{\gamma}(\mathcal{N}(v))\in\mathbb{R}^{d_{out}}$ is the raw neighborhood alignment difficulty of each node, where $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T})$. $\mathcal{A}$ is the adjacency matrix of their respective graphs. We further combine the aggregated difficulties with representations $\mathbf{h}_{v}$, where $\mathbf{h}_{v}$ serve as residual information to strengthen the connections between difficulties and the corresponding representations, as follows.
(11) $\gamma(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Concat}(\hat{\gamma}(\mathcal{N}(v)),\mathbf{h}_{v})\bigg{)}$
Where $\gamma(\mathcal{N}(v))\in\mathbb{R}^{d_{out}}$ is the neighborhood alignment difficulty for each node $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T})$.
Inter-graph Consideration and Certainty Computation.
A match connects two nodes within different graphs, each node can have a distinct neighborhood alignment difficulty. In this step, Gen considers such inter-graph relationships to evaluate the certainty for each match, balancing the neighborhood alignment difficulties of the involved nodes. Specifically, it concatenates the neighborhood alignment difficulties and leverages a MLP to compute the matching certainty, approximating $\mathbf{1}_{s,t}$ as follows.
(12) $\mathbf{1}_{s,t}=\text{MLP}\bigg{(}\text{Concat}(\gamma(s),\gamma(t))\bigg{)}$
Where $\mathbf{1}_{s,t}\in\mathbb{R}^{d_{out}}$ represents the certainty of each match, enabling more informed decisions by considering the interdependencies between matches.
4.4. Prediction and Training Objective
Predictions. Based on the certainty of each match, Gen softly selects optimal matches in a single step for each node, and the matching cost for each node is calculated as follows.
(13) $c(s)=\sum_{i\in\mathcal{V}_{T}}\mathbf{1}_{s,i}\cdot c(s,i)\quad\text{and}\quad c(t)=\sum_{i\in\mathcal{V}_{s}}\mathbf{1}_{i,t}\cdot c(i,t)$
Finally, the total matching cost of each graph is calculated as the sum of certainty-weighted node matching costs, according to Eq. 5. Then, the GED is computed based on the average of the overall matching costs of the input graph pair.
(14) $\begin{gathered}c(S,T)_{S}=\sum_{s\in\mathcal{V}_{S}}c(s)\quad\text{and}\quad c(S,T)_{T}=\sum_{t\in\mathcal{V}_{S}}c(t)\\ \text{GED}(S,T)=\text{Linear}\bigg{(}\text{Mean}(c(S),c(T))\bigg{)}\end{gathered}$
Where $c(S,T)_{S}$ and $c(S,T)_{T}$ are the total costs of mappings established for $S$ and $T$, respectively. $\text{GED}(S,T)\in\mathbb{R}$ is the predicted GED, $\text{Linear}(\cdot)$ reduces the dimensionality from $d$ to $1$. This process indirectly encourages consistent mappings between the $S$ and $T$, or the average matching cost increases accordingly.
Training Objective. Given the GED values as supervision signals, which span from 0 to potentially large numbers, we adopt Huber loss (Huber, 1964) for Gen and baselines to enable stable training. It combines the properties of Mean Squared Error (MSE) and Mean Absolute Error (MAE) losses to handle both small and large errors. Given the predicted GED $\hat{y}$ and the ground-truth GED $y$, the loss is computed as follows.
(15) $\mathcal{L}(\hat{y},y)=\begin{cases}\frac{1}{2}(y-\hat{y})^{2}&\text{for }|y-\hat{y}|\leq\delta,\\ \delta(|y-\hat{y}|-\frac{1}{2}\delta)&\text{otherwise. }\end{cases}$
Here, $\delta$ is the threshold for the loss function transition. For errors smaller than $\delta$, it behaves like MSE loss, offering a smooth gradient. For larger errors, it switches to MAE loss, preventing large gradients and improving training stability.
4.5. Complexity Analysis
The complexity of GNN Encoder is $\mathcal{O}(k\cdot|\mathcal{E}|\cdot d_{hidden})$, where $k$ is the number of layers, $\mathcal{E}$ is the number of edges in the input graph, $d_{hidden}$ is the hidden dimensionality during message-passing, we further reduce it to $d_{out}$, and the generated node representations can be reused for each graph as they are pair-independent. The complexity of the Node Matching Cost Estimation is $\mathcal{O}(\text{Max}(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}\cdot d_{out})$, where $|\mathcal{V}_{S}|$ and $|\mathcal{V}_{T}|$ are the numbers of nodes in the source and target graphs, respectively. The complexities of Proactive Optimization Strategy is $\mathcal{O}(2\cdot(\text{Max}(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}\cdot d_{out}+|\mathcal{E}|\cdot d_{out}))$.
5. Evaluations
In this section, we evaluate Gen against twelve baselines, including two conventional inexact algorithms, on three commonly used GED datasets and synthetic graphs, focusing on: (1) Effectiveness: How accurate is Gen compared with baselines in the GED computation and graph retrieval under a uniform cost setting; (2) Time Efficiency: How efficient is Gen during training and inference; (3) Adaptation to Varying Cost Settings: Can Gen and baselines accommodate non-uniform and varying cost settings; (4) Ablation Study: How effective are our proposed strategy; (5) Hyperparameter Senstitivity: How sensitive is Gen to the number of layers, dimensionality, and GNN backbones; (6) Performance on Larger Synthetic Graphs: How do the models perform on larger synthetic graphs.
5.1. Experimental Setup
5.1.1. Dataset
We conduct experiments on three commonly used GED datasets: AIDS700nef, Linux and IMDBMulti. The graphs within these datasets and the GED values between graph pairs under a uniform cost setting are provided by Bai et al. (2019), resulting in a rich set of graph pairs. The exact GED is computed with A* (Abu-Aisheh et al., 2015) for AIDS700nef and LINUX. Since the exact GED computation is intractable for IMDBMulti, the smallest distances computed by inexact algorithms such as Beam (Neuhaus et al., 2006), VJ (Fankhauser et al., 2011), and Hungarian (Riesen and Bunke, 2009) are adopted as the training targets. For AIDS700nef, we compute GEDs under two non-uniform settings using A* (Abu-Aisheh et al., 2015). Additionally, we generate synthetic graphs, among Erdős-Rényi, power-law, and Watts–Strogatz small-world graphs with a node size ranging from 100 to 400, as extreme data to to assess model performance on larger graphs. Each synthetic graph’s target graph is created by applying $\Delta$ distinct operations, with the number of operations ranging from 10-30% of the edge count. Following prior work (Li et al., 2019; Bai and Zhao, 2021; Piao et al., 2023), $\Delta$ serves as the ground-truth GED for synthetic pairs. We generate 51,200 graphs for training and another 38,400 graphs in three node size groups: $[100,200]$, $[200,300]$, and $[300,400]$, for testing. The statistics of graphs can be found in Table 1.
Table 1. Statistics of datasets.
Dataset #Graphs #Pairs Avg. $|\mathcal{V}|$ Avg. $|\mathcal{E}|$ Max. $|\mathcal{V}|$ Max. $|\mathcal{E}|$
AIDS700nef 700 490K 8.90 8.80 10 14
LINUX 1000 1M 7.58 6.94 10 13
IMDBMulti 1500 2.25M 13.00 65.94 89 1467
Synthetic 51,200 25,6 K 230.24 1789.02 400 8552
5.1.2. Baselines
We includes representative and state-of-the-art GED computation models as baselines: (1) Graph Comparison Methods: GMN-emb (Li et al., 2019), TaGSim (Bai and Zhao, 2021), Greed (Ranjan et al., 2022) and Eric (Zhuo and Tan, 2022); (2) Node Comparison Method: GMN-match (Li et al., 2019), SimGNN (Bai et al., 2019), GraphSim (Bai et al., 2020), GOTSim (Doan et al., 2021), GEDGNN (Piao et al., 2023), GraphEdX (Jain et al., 2024). Additionally, we compare with well-known approximate algorithms, such as the Volgenant–Jonker (VJ) algorithm (Fankhauser et al., 2011) and Beam (Neuhaus et al., 2006). The VJ algorithm is a linear sum assignment method that greedily performs bipartite graph matching to minimize costs, while Beam employs a best-first strategy, expanding partial solutions with the lowest-cost matches at each step to approximate GED. Hungarian (Riesen and Bunke, 2009) is not included as it has been incorporated in GOTSim. We generally use the official implementations and hyperparameters provided by the authors, with the VJ and Beam implementations sourced from (Yang and Zou, 2021). For models predicting the normalized GED value, we remove their last sigmoid layer to ensure a fair comparison.
5.1.3. Experimental Protocol
We partitioned the datasets into training and test graphs at a 4:1 ratio, with 20% of the training graphs used as the validation graphs. Training graph pairs are derived from the training graphs. All models are trained on a NVIDIA A6000 GPU with AMD EPYC 7282 CPU. The batch size is set to 128 graph pairs for training, 2048 for validation and test. Each epoch comprises 100 iterations due to the abundance of graph pairs, except for synthetic graphs. Each model is trained for ten epochs as a warm-up phase and tested on the validation set every epoch. Early stopping is applied with a patience of 50, meaning training stops if the validation loss fails to decrease for 50 consecutive validation steps to prevent overfitting, while ensuring that each model is sufficiently trained.
5.1.4. Implementation Details
We apply a 5-layer Graph Isomorphism Network (GIN) (Xu et al., 2019) with jump connections as GNN encoder to neighborhood information, whose hidden state and output are of 64 dimensionality. All MLPs in this work have adopted layer normalization and applied ReLU as the activation function. The outputs of these MLPs are of a dimensionality of 32. We used Adam as the optimizer, with a fixed learning rate of 0.001.
5.1.5. Evaluation Metrics
We evaluate models based on prediction error and ranking ability. Prediction error measures the point-wise difference between predicted and ground-truth GED values. Ranking ability is assessed by simulating retrieval scenarios in graph databases, where training and validation graphs serve as data graphs and test graphs as queries. During testing, each model calculates the GED between the query and data graphs, returning the most relevant data graphs with the lowest predicted GED for each query. For prediction error, we use Root Mean Square Error (RMSE) and Mean Absolute Error (MAE). Similar to MSE, RMSE is sensitive to outliers but presents errors in the original units for better interpretability, while MAE captures general errors. For ranking performance, we use Spearman’s Rank Correlation Coefficient ($\rho$), Kendall’s Rank Correlation Coefficient ($\tau$), and Precision at $k$ ($p@k$). $\rho$ sensitive to the magnitude of the differences in ranks, while $\tau$ evaluates relative order of pairs. $p@k$ calculates the intersection between the predicted top $k$ data graphs and the ground-truth top $k$ results. We sort all data graphs for each query based on both the ground truth GEDs and predictions in ascending order and compare the rankings. We report the average and standard deviation of results across five runs.
Table 2. Effectiveness of models on test sets. We bold the best and underline the second models. ’OOM’ indicates out of memory.
Methods AIDS700nef LINUX IMDBMulti
RMSE $\downarrow$ MAE $\downarrow$ $\rho$uparrow $\tau$uparrow p@10 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $\rho$uparrow $\tau$uparrow p@10 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $\rho$uparrow $\tau$uparrow p@10 $\uparrow$
VJ $9.278_{\pm\text{0.000}}$ $8.709_{\pm\text{0.000}}$ $0.422_{\pm\text{0.000}}$ $0.296_{\pm\text{0.000}}$ $0.269_{\pm\text{0.000}}$ $3.407_{\pm\text{0.000}}$ $2.692_{\pm\text{0.000}}$ $0.728_{\pm\text{0.000}}$ $0.574_{\pm\text{0.000}}$ $0.934_{\pm\text{0.000}}$
Beam $5.111_{\pm\text{0.000}}$ $4.380_{\pm\text{0.000}}$ $0.486_{\pm\text{0.000}}$ $0.345_{\pm\text{0.000}}$ $0.371_{\pm\text{0.000}}$ $3.315_{\pm\text{0.000}}$ $2.841_{\pm\text{0.000}}$ $0.734_{\pm\text{0.000}}$ $0.575_{\pm\text{0.000}}$ $0.445_{\pm\text{0.000}}$
SimGNN $1.036_{\pm\text{0.040}}$ $1.075_{\pm\text{0.084}}$ $0.847_{\pm\text{0.008}}$ $0.670_{\pm\text{0.008}}$ $0.897_{\pm\text{0.005}}$ $0.621_{\pm\text{0.076}}$ $0.434_{\pm\text{0.066}}$ $0.919_{\pm\text{0.009}}$ $0.760_{\pm\text{0.012}}$ $0.869_{\pm\text{0.065 }}$ $13.721_{\pm\text{2.163}}$ $6.863_{\pm\text{1.161}}$ $0.899_{\pm\text{0.006}}$ $0.781_{\pm\text{0.006}}$ $0.978_{\pm\text{0.001}}$
GMN-match $1.068_{\pm\text{0.032}}$ $0.811_{\pm\text{0.021}}$ $0.873_{\pm\text{0.006}}$ $0.701_{\pm\text{0.008}}$ $0.899_{\pm\text{0.004}}$ $\underline{0.267_{\pm\text{0.007}}}$ $0.126_{\pm\text{0.008}}$ $\underline{0.946_{\pm\text{0.001}}}$ $0.805_{\pm\text{0.002}}$ $0.967_{\pm\text{0.009}}$ $\underline{8.401_{\pm\text{2.061}}}$ $\underline{3.694_{\pm\text{0.925}}}$ $\textbf{0.938}_{\pm\textbf{0.007}}$ $\textbf{0.843}_{\pm\textbf{0.017}}$ $\underline{0.985_{\pm\text{0.004}}}$
GMN-emb $1.645_{\pm\text{0.081}}$ $1.145_{\pm\text{0.033}}$ $0.854_{\pm\text{0.006}}$ $0.679_{\pm\text{0.006}}$ $0.793_{\pm\text{0.025}}$ $0.334_{\pm\text{0.033}}$ $0.174_{\pm\text{0.010}}$ $0.944_{\pm\text{0.001}}$ $0.803_{\pm\text{0.002}}$ $0.974_{\pm\text{0.014}}$ $12.340_{\pm\text{1.403}}$ $5.296_{\pm\text{0.707}}$ $0.926_{\pm\text{0.005}}$ $0.820_{\pm\text{0.010}}$ $0.966_{\pm\text{0.034}}$
GraphSim $2.017_{\pm\text{0.793}}$ $1.519_{\pm\text{0.580}}$ $0.382_{\pm\text{0.381}}$ $0.298_{\pm\text{0.306}}$ $0.365_{\pm\text{0.447}}$ $2.283_{\pm\text{1.008}}$ $1.721_{\pm\text{0.800}}$ $0.206_{\pm\text{0.371}}$ $0.175_{\pm\text{0.319}}$ $0.276_{\pm\text{0.359}}$ $159.571_{\pm\text{0.012}}$ $70.917_{\pm\text{0.002}}$ $0.003_{\pm\text{0.000}}$ $0.002_{\pm\text{0.000}}$ $0.001_{\pm\text{0.000}}$
GOTSim $1.584_{\pm\text{0.347}}$ $1.198_{\pm\text{0.249}}$ $0.691_{\pm\text{0.183}}$ $0.525_{\pm\text{0.149}}$ $0.783_{\pm\text{0.058}}$ $1.560_{\pm\text{0.621}}$ $1.242_{\pm\text{0.460}}$ $0.666_{\pm\text{0.363}}$ $0.528_{\pm\text{0.281}}$ $0.550_{\pm\text{0.066}}$ $129.623_{\pm\text{13.795}}$ $52.248_{\pm\text{7.199}}$ $0.611_{\pm\text{0.250}}$ $0.492_{\pm\text{0.208}}$ $0.474_{\pm\text{0.201}}$
TaGSim $1.113_{\pm\text{0.058}}$ $0.844_{\pm\text{0.039}}$ $0.841_{\pm\text{0.012}}$ $0.663_{\pm\text{0.014}}$ $0.896_{\pm\text{0.009}}$ $0.678_{\pm\text{0.066}}$ $0.438_{\pm\text{0.051}}$ $0.913_{\pm\text{0.010}}$ $0.758_{\pm\text{0.012}}$ $0.909_{\pm\text{0.029}}$ $11.565_{\pm\text{3.310}}$ $5.451_{\pm\text{1.340}}$ $0.873_{\pm\text{0.053}}$ $0.764_{\pm\text{0.059}}$ $0.979_{\pm\text{0.008}}$
ERIC $1.259_{\pm\text{0.316}}$ $0.971_{\pm\text{0.228}}$ $0.782_{\pm\text{0.097}}$ $0.605_{\pm\text{0.095}}$ $0.836_{\pm\text{0.111}}$ $0.271_{\pm\text{0.130}}$ $\underline{0.120_{\pm\text{0.104}}}$ $0.944_{\pm\text{0.007}}$ $0.803_{\pm\text{0.012}}$ $0.973_{\pm\text{0.027}}$ $9.866_{\pm\text{0.662}}$ $4.462_{\pm\text{0.292}}$ $0.932_{\pm\text{0.005}}$ $0.834_{\pm\text{0.008}}$ $0.977_{\pm\text{0.006}}$
Greed $0.900_{\pm\text{0.016}}$ $0.705_{\pm\text{0.012}}$ $0.885_{\pm\text{0.001}}$ $0.715_{\pm\text{0.002}}$ $0.930_{\pm\text{0.005}}$ $0.441_{\pm\text{0.003}}$ $0.327_{\pm\text{0.002}}$ $0.939_{\pm\text{0.001}}$ $0.791_{\pm\text{0.001}}$ $0.960_{\pm\text{0.002}}$ $19.974_{\pm\text{20.988}}$ $6.948_{\pm\text{5.880}}$ $0.890_{\pm\text{0.039}}$ $0.778_{\pm\text{0.047}}$ $0.837_{\pm\text{0.303}}$
GEDGNN $1.607_{\pm\text{0.618}}$ $1.148_{\pm\text{0.354}}$ $0.747_{\pm\text{0.114}}$ $0.577_{\pm\text{0.104}}$ $0.739_{\pm\text{0.155}}$ $0.313_{\pm\text{0.098}}$ $0.188_{\pm\text{0.071}}$ $0.945_{\pm\text{0.005}}$ $\underline{0.806_{\pm\text{0.007}}}$ $\underline{0.993_{\pm\text{0.005}}}$ $15.743_{\pm\text{9.533}}$ $7.070_{\pm\text{3.471}}$ $0.897_{\pm\text{0.036}}$ $0.785_{\pm\text{0.046}}$ $0.934_{\pm\text{0.076}}$
GraphEdX $\textbf{0.735}_{\pm\textbf{0.015}}$ $\underline{0.581_{\pm\text{0.011}}}$ $\underline{0.910_{\pm\text{0.003}}}$ $\underline{0.749_{\pm\text{0.004}}}$ $\underline{0.951_{\pm\text{0.003}}}$ $0.378_{\pm\text{0.135}}$ $0.222_{\pm\text{0.081}}$ $0.941_{\pm\text{0.006}}$ $0.797_{\pm\text{0.011}}$ $0.973_{\pm\text{0.021}}$ OOM OOM OOM OOM OOM
Ours (Gen) $\underline{0.738_{\pm\text{0.007}}}$ $\textbf{0.577}_{\pm\textbf{0.005}}$ $\textbf{0.915}_{\pm\textbf{0.002}}$ $\textbf{0.756}_{\pm\textbf{0.003}}$ $\textbf{0.963}_{\pm\textbf{0.003}}$ $\textbf{0.166}_{\pm\textbf{0.009}}$ $\textbf{0.084}_{\pm\textbf{0.008}}$ $\textbf{0.948}_{\pm\textbf{0.000}}$ $\textbf{0.810}_{\pm\textbf{0.001}}$ $\textbf{0.995}_{\pm\textbf{0.004}}$ $\textbf{7.531}_{\pm\textbf{0.293}}$ $\textbf{3.492}_{\pm\textbf{0.128}}$ $\underline{0.937_{\pm\text{0.002}}}$ $\underline{0.839_{\pm\text{0.002}}}$ $\textbf{0.987}_{\pm\textbf{0.001}}$
5.2. Results
5.2.1. Effectiveness
In this experiment, we compared the performance of our proposed Gen against all baselines under a uniform cost setting (operation costs set to 1). Results for the VJ and Beam algorithms on IMDBMulti were excluded, as IMDBMulti’s training targets are the smallest distances calculated by these algorithms. As shown in Table 2, Gen consistently ranks first or second across all metrics on the three datasets, demonstrating its competitive performance. GNN-based methods generally outperform conventional ones, as the former leverage enriched information extracted by GNNs. Graph comparison methods, such as Greed and Eric, outperform most node comparison competitors. This suggests that there is no clear winner between the two categories in this setting, where all models predict inherent graph differences. Most node comparison methods underperformed due to their reliance on node-level guidance, which fails to evaluate matches from a graph perspective. Among these methods, GraphEdX demonstrated advantages on the node-labeled AIDS700nef dataset by using XOR operation as a node distance measure. However, it underperformed on the LINUX dataset because it is unlabeled. Additionally, it encountered out-of-memory errors on IMDBMulti with larger graphs because it, by design, requires padding all graphs to the same size, whereas other models, such as GEDGNN and Gen, pad the graphs within graph pairs instead. GMN-match iteratively fuses and propagates node representations across graphs. It showed competitive performance on two datasets but suffered from error accumulation due to its greedy node-level fusion. Similarly, GraphSim’s BFS-based matrix reordering led to instability, as indicated by large standard deviations. In contrast to these node comparison methods, Gen introduces a proactive optimization strategy that leverages node alignment difficulty as guidance and optimizes it by considering interdependencies between matches. This strategy enables Gen to achieve up to a 37.8% error reduction in RMSE and a 30% reduction in MAE on the LINUX dataset. Additionally, Gen generally improves ranking performance by up to 4.1%, with a marginal decrease of up to 0.4% in Kendall’s tau $\tau$ on the IMDBMulti dataset. This suggests potential optimization for datasets like IMDBMulti, whose graphs have a low diameter of up to 2. A case study comparing the ranking results of Gen against the ground truth provided by A* is available in Figure 3.
Figure 3. Case Study of ranking results on AIDS700nef (labeled) and IMDBMulti (unlabeled).
Table 3. We add ’Rank’ to indicate the current ranking based on RMSE and show the improvements over the ranking for the AIDS700nef dataset in Table 2 under a uniform cost setting, as shown in parentheses.
Dataset Setting 1 Setting 2
[…]
SimGNN $1.923_{\pm\text{0.082}}$ $1.487_{\pm\text{0.064}}$ $0.859_{\pm\text{0.010}}$ $0.680_{\pm\text{0.013}}$ 5 (-1) $1.100_{\pm\text{0.024}}$ $0.862_{\pm\text{0.015}}$ $0.955_{\pm\text{0.002}}$ $0.820_{\pm\text{0.003}}$ 4 (0)
GMN-match $4.874_{\pm\text{0.054}}$ $3.421_{\pm\text{0.054}}$ $0.423_{\pm\text{0.017}}$ $0.295_{\pm\text{0.014}}$ 11 (-6) $4.761_{\pm\text{0.152}}$ $3.317_{\pm\text{0.086}}$ $0.471_{\pm\text{0.009}}$ $0.344_{\pm\text{0.009}}$ 11 (-6)
GMN-emb $5.742_{\pm\text{0.309}}$ $4.070_{\pm\text{0.110}}$ $0.448_{\pm\text{0.009}}$ $0.311_{\pm\text{0.006}}$ 13 (-3) $8.694_{\pm\text{0.117}}$ $6.093_{\pm\text{0.097}}$ $0.182_{\pm\text{0.028}}$ $0.125_{\pm\text{0.020}}$ 13 (-3)
GraphSim $3.874_{\pm\text{1.636}}$ $2.707_{\pm\text{1.047}}$ $0.356_{\pm\text{0.419}}$ $0.283_{\pm\text{0.336}}$ 9 (+2) $3.226_{\pm\text{1.574}}$ $2.488_{\pm\text{1.212}}$ $0.380_{\pm\text{0.458}}$ $0.320_{\pm\text{0.388}}$ 9 (+2)
GOTSim $3.316_{\pm\text{0.211}}$ $2.635_{\pm\text{0.186}}$ $0.768_{\pm\text{0.058}}$ $0.577_{\pm\text{0.053}}$ 8 (0) $2.754_{\pm\text{0.049}}$ $2.173_{\pm\text{0.034}}$ $0.760_{\pm\text{0.021}}$ $0.564_{\pm\text{0.017}}$ 8 (0)
TaGSim $2.120_{\pm\text{0.051}}$ $1.605_{\pm\text{0.017}}$ $0.844_{\pm\text{0.003}}$ $0.662_{\pm\text{0.003}}$ 6 (0) $1.188_{\pm\text{0.081}}$ $0.864_{\pm\text{0.032}}$ $0.950{\pm\text{0.004}}$ $0.816_{\pm\text{0.005}}$ 6 (0)
ERIC $2.319_{\pm\text{0.173}}$ $1.826_{\pm\text{0.116}}$ $0.784_{\pm\text{0.035}}$ $0.599_{\pm\text{0.033}}$ 7 (0) $1.587_{\pm\text{0.138}}$ $1.250_{\pm\text{0.096}}$ $0.911_{\pm\text{0.011}}$ $0.747_{\pm\text{0.014}}$ 7 (0)
Greed $4.607_{\pm\text{0.045}}$ $3.396_{\pm\text{0.040}}$ $0.494_{\pm\text{0.012}}$ $0.348_{\pm\text{0.009}}$ 10 (-7) $5.874_{\pm\text{0.072}}$ $4.200_{\pm\text{0.060}}$ $0.237_{\pm\text{0.013}}$ $0.170_{\pm\text{0.010}}$ 12 (-9)
GEDGNN $1.782_{\pm\text{0.921}}$ $1.331_{\pm\text{0.598}}$ $0.916_{\pm\text{0.082}}$ $0.770_{\pm\text{0.104}}$ 4 (+5) $1.102_{\pm\text{0.630}}$ $0.850_{\pm\text{0.466}}$ $0.953_{\pm\text{0.049}}$ $0.830_{\pm\text{0.080}}$ 5 (+4)
GraphEdX $1.249_{\pm\text{0.014}}$ $0.978_{\pm\text{0.008}}$ $0.934_{\pm\text{0.002}}$ $0.786_{\pm\text{0.002}}$ 3 (-2) $0.714_{\pm\text{0.012}}$ $0.567_{\pm\text{0.009}}$ $0.977_{\pm\text{0.001}}$ $0.871_{\pm\text{0.002}}$ 3 (-2)
Ours (Gen) $\underline{1.061_{\pm\text{0.032}}}$ $\underline{0.771_{\pm\text{0.033}}}$ $\underline{0.952_{\pm\text{0.003}}}$ $\underline{0.823_{\pm\text{0.006}}}$ 2 (0) $\underline{0.545_{\pm\text{0.007}}}$ $\underline{0.409_{\pm\text{0.007}}}$ $\underline{0.984_{\pm\text{0.000}}}$ $\underline{0.893_{\pm\text{0.001}}}$ 2 (0)
TaGSim mix $7.106_{\pm\text{0.137}}$ $5.249_{\pm\text{0.159}}$ $0.307_{\pm\text{0.024}}$ $0.217_{\pm\text{0.017}}$ 14 (new) $4.088_{\pm\text{0.171}}$ $3.087_{\pm\text{0.182}}$ $0.510_{\pm\text{0.014}}$ $0.382_{\pm\text{0.011}}$ 10 (new)
GraphEdX mix $5.024_{\pm\text{0.411}}$ $4.081_{\pm\text{0.357}}$ $0.620_{\pm\text{0.058}}$ $0.446_{\pm\text{0.047}}$ 12 (new) $12.918_{\pm\text{1.144}}$ $10.919_{\pm\text{1.334}}$ $0.218_{\pm\textbf{0.051}}$ $0.142_{\pm\textbf{0.038}}$ 14 (new)
Ours (Gen mix) $\textbf{0.898}_{\pm\textbf{0.044}}$ $\textbf{0.642}_{\pm\textbf{0.043}}$ $\textbf{0.979}_{\pm\textbf{0.002}}$ $\textbf{0.881}_{\pm\textbf{0.006}}$ 1 (new) $\textbf{0.479}_{\pm\textbf{0.013}}$ $\textbf{0.353}_{\pm\textbf{0.015}}$ $\textbf{0.988}_{\pm\textbf{0.001}}$ $\textbf{0.905}_{\pm\textbf{0.003}}$ 1 (new)
(a) Training Time.
(b) # Epochs.
(c) Inference Time.
Figure 4. Time Efficiency of GSC models. The bars with black edges represent node comparison-based methods.
5.2.2. Time Efficiency
In this experiment, we evaluated the training and inference times of models. Results for VJ and Beam were omitted, as prior experiments showed learning-based methods outperform them in efficiency (Bai et al., 2019; Zhang et al., 2021). Figures 4(a) and 4(b) illustrate the average time and epochs needed for convergence. Our proposed Gen ranks fourth or fifth fastest in training time, outperforming all node comparison (NC) models, and competing with the slowest graph comparison (GC) model, Eric, which constrains node-graph alignment during training. Although SimGNN shares a similar batch runtime with ours, it requires more epochs to converge. Figure 4(c) illustrates the total inference time on testset, GC methods are generally more time-efficient than NC ones, all of latter exhibit at least quadratic time complexity. Compared with the fastest NC method, SimGNN, Gen incurs an 11.2% increase in inference time but achieves up to 80.6% error reduction. GEDGNN and GraphEdX are among the slowest, as they establish mappings in a matrix formulation based on matching costs and require 20 refinement iterations to ensure performance. Similarly, GMN-match requires up to 5 iterative refinements for comparable accuracy. These refinements significantly increase computational demand as the number and size of graph pairs grow, highlighting by their increased time consumption on the LINUX and IMDBMulti datasets. Compared with the most relevant and competitive GMN-match, Gen reduces inference time by up to 72.7%. This improvement comes from the effective graph perspective guidance provided our proposed strategy, which enable Gen to make more informed decisions, minimizing the need for refinements, and leading to faster training and inference.
5.2.3. Adaptation to Varying Costs
In this experiment, we assess the adaptability of models to non-uniform and varying cost settings using the node-labeled AIDS700nef dataset. We evaluate two cost settings: Setting 1 assigns costs of 1 for node substitution, 2 for node deletion, 1 for node insertion, 3 for edge deletion, and 1 for edge insertion, denoted as [1, 2, 1, 3, 1]. In Setting 2, the corresponding costs are [3, 2, 3, 0, 2]. The results, presented in Table 3, show that: (i) Under non-uniform cost settings: Our proposed Gen model demonstrates superior adaptability, achieving the best performance across all metrics for both settings. The performance of Greed degraded sharply, likely due to its use of symmetric Euclidean distance, whereas under these settings, the matching cost from $s$ to $t$ may differ from that of $t$ to $s$, making it asymmetric. GMN-match also underperformed, as it greedily fuses nodes based on inherent node similarity, which can lead to significant error accumulations. Notably, all ’Rank’ improvements occurred for node comparison methods, suggesting that they are more suitable for flexible GED computation. (ii) Under varying costs: We trained Gen mix, TaGSim mix, and GraphEdX mix, these consider operation types and costs, using a mixed dataset under two cost settings. Results show that Gen mix substantially outperforms all models trained under a single cost setting and TaGSim mix and GraphEdX mix trained under varying settings, demonstrating better adaptability. The sharp performance degradation of TaGSim mix and GraphEdX mix highlights their key limitation, as they treat operation costs as irrelevant when determining the mappings, resulting in static mappings that are suboptimal under non-uniform settings. In contrast, the improvement of Gen mix over Gen demonstrates that our principled solution enhances the model’s understanding of GED computation. By preserving consistent node distances and dynamically adjusting them based on specific operation costs, it enables better adaptation to varying cost settings.
Table 4. Ablations Study on our proposed components.
Dataset AIDS700nef LINUX IMDBMulti
RMSE $\downarrow$ p@10 $\uparrow$ RMSE $\downarrow$ p@10 $\uparrow$ RMSE $\downarrow$ p@10 $\uparrow$
Gen wo global $0.808_{\pm\text{0.038}}$ $0.954_{\pm\text{0.009}}$ $0.178_{\pm\text{0.013}}$ $0.993_{\pm\text{0.005}}$ $9.415_{\pm\text{0.977}}$ $0.972_{\pm\text{0.023}}$
Gen wo interdependencies $1.101_{\pm\text{0.041}}$ $0.905_{\pm\text{0.006}}$ $0.296_{\pm\text{0.034}}$ $\underline{0.990_{\pm\text{0.005}}}$ $14.391_{\pm\text{2.568}}$ $0.982_{\pm\text{0.004}}$
Gen wo intra-graph $\underline{0.746_{\pm\text{0.011}}}$ $\underline{0.958_{\pm\text{0.002}}}$ $\textbf{0.159}_{\pm\textbf{0.015}}$ $\textbf{0.995}_{\pm\textbf{0.003}}$ $\underline{8.439_{\pm\text{0.656}}}$ $\underline{0.984_{\pm\text{0.005}}}$
Gen $\textbf{0.738}_{\pm\textbf{0.007}}$ $\textbf{0.963}_{\pm\textbf{0.003}}$ $\underline{0.166_{\pm\text{0.009}}}$ $\textbf{0.995}_{\pm\textbf{0.004}}$ $\textbf{7.531}_{\pm\textbf{0.293}}$ $\textbf{0.987}_{\pm\textbf{0.001}}$
5.2.4. Ablation Study
In this experiment, we evaluate the effectiveness of our proposed strategy. We test three variants: (1) Gen wo global: This variant replaces the global cost ratio with absolute node matching costs to estimate alignment difficulties. (2) Gen wo interdependencies: This variant reweighs matches solely based on their global cost ratios, where higher ratios result in lower weights. (3) Gen wo intra-graph: This variant does not compute the neighborhood alignment difficulty and considers inter-graph interdependency based only on node alignment difficulties. Note that the inter-graph interdependency cannot be ablated separately as it produces the decision for each match. As shown in Table 4, Gen wo interdependencies greedily reweighs matches based on global cost ratios, leading to a significant performance drop across all datasets, with RMSE increasing by up to 91.1%. On the other hand, Gen wo global incurs a smaller increase, at most 25%, indicating that considering interdependencies is more crucial than the global cost ratio. Meanwhile, ignoring the difficulties of neighboring nodes has a smallest impact. Gen wo intra-graph showcased comparable performance on AIDS700nef and LINUX, yet incurred a 12.1% increase in RMSE on IMDBMulti. This is likely because the matches in IMDBMulti are more ambiguous, making the neighborhood alignment difficulty a crucial factor.
Figure 5. Hyperparameter sensitivity.
5.2.5. Hyperparameter Sensitivity
We analyze Gen’s sensitivity to key hyperparameters such as GNN backbones, number of layers, and dimensionality. As shown in Figure 5, GIN proves to be the optimal backbone for Gen, consistently achieving the lowest RMSE. Adding layers generally improves performance, with the effect varying across datasets. On AIDS700nef and LINUX, performance increases steadily with more layers, while on IMDBMulti, it peaks at four layers for RMSE reduction. Similarly, increasing dimensionality typically reduces RMSE, but on IMDBMulti, performance worsens beyond 32/16 dimensions, likely due to overfitting.
Figure 6. Performance of models on larger synthetic graphs across diverse node size groups.
5.2.6. Performance on Larger Synthetic Graphs
GED is typically applied to small-sized graphs, as its exact computation becomes challenging for graphs with more than 16 nodes (Blumenthal and Gamper, 2020). To evaluate the performance of GSC models under extreme conditions, we followed (Piao et al., 2023) and assessed the performance of GSC models on larger synthetic graphs. We compared GED predictions among baseline methods and our proposed Gen against ground-truth GEDs. Figure 6 illustrates the results for the best node and graph comparison models during this experiment, GEDGNN and TaGSim, against our proposed Gen. The results show that Gen achieved the most accurate predictions, with data points closely clustered around the line representing perfect alignment between predictions and GEDs across all node size groups. In contrast, GEDGNN often underestimated, while TaGSim diverged more from the line and produced higher outlier predictions as node size grows.
6. Conclusion
This paper presents Graph Edit Networks (Gen), a learning-based model that captures the inherent flexibility of graph edit distance (GED) computation with an efficient proactive optimization strategy. We highlight a key limitation in existing methods in handling flexible GED computation: they neglect or treat operation costs as irrelevant when determining the mappings, resulting in static mappings across cost settings. To address this issue, we follow the GED computation principle and propose incorporating operation costs before establishing mappings. Instead of reactively refining mappings with insufficient node-level guidance, we introduce a novel proactive optimization strategy. Coupled with a difficulty propagation mechanism, this strategy enables effective guidance from a graph perspective, allowing match selections in a single step. Experiments on both real-world and synthetic datasets validate the model’s effectiveness, efficiency, and adaptability. Our proposed approach can be applied in graph retrieval scenarios, where Gen can return the top $k$ relevant graphs based on a query or rerank candidate graphs generated by fast graph comparison models to improve precision.