### v2

# Flexible Graph Similarity Computation With A Proactive Optimization Strategy

Abstract
Graph Edit Distance (GED) offers a principled and flexible measure of graph similarity, as it quantifies the minimum cost needed to transform one graph into another with customizable edit operation costs. Despite recent learning-based efforts to approximate GED via vector space representations, existing methods struggle with adapting to varying operation costs. Furthermore, they suffer from inefficient, reactive mapping refinements due to reliance on isolated node-level distance as guidance. To address these issues, we propose Gen, a novel learning-based approach for flexible GED approximation. Gen addresses the varying costs adaptation by integrating operation costs prior to match establishment, enabling mappings to dynamically adapt to cost variations. Furthermore, Gen introduces a proactive guidance optimization strategy that captures graph-level dependencies between matches, allowing informed matching decisions in a single step without costly iterative refinements. Extensive evaluations on real-world and synthetic datasets demonstrate that Gen achieves up to 37.8% reduction in GED approximation error and 72.7% reduction in inference time compared with state-of-the-art methods, while consistently maintaining robustness under diverse cost settings and graph sizes.
1 Introduction
Measuring graph similarity is a fundamental problem in graph mining, with wide-ranging applications in retrieval, classification, and recommendation tasks [8, 11, 18, 27, 32, 33, 34]. Among various similarity measures, Graph Edit Distance (GED) is particularly appealing due to its flexibility: it defines the similarity between two graphs as the minimum cost of transforming one into another through node/edge insertion, deletion, or substitution, under customizable cost settings. This flexibility allows GED to be tailored to specific applications. For example, in graph databases for recommender systems, higher costs can be assigned to nodes/edges representing key user preferences. This prioritizes the retrieval of graphs that preserve these preferences, enhancing personalized recommendations.
Traditionally, exact GED computation relies on A*-like solvers [2, 8, 30], which sequentially construct node/edge mappings. In this procedure, operation costs critically influence matching decisions, resulting in varying optimal mappings under different cost settings, as illustrated in Figure 1. Moreover, the sequential manner maintains structural consistency with prior decisions, and inherently models dependencies among node matches induced by intra-graph connections and inter-graph discrepancies. Such a procedure effectively embodies two fundamental principles of GED computation: the influence of operation costs and the dependencies between node matches. However, exhaustive search incurs prohibitive computational costs. Even for graphs with as few as 16 nodes, exact GED computation remains infeasible within practical time limits [6]. This bottleneck severely restricts the applicability of traditional methods in time-sensitive graph retrieval scenarios.
In response, recent advances leverage Graph Neural Networks (GNNs) [7, 15, 26, 28] to approximate GED by embedding graphs into vector spaces [3, 4, 5, 9, 13, 14, 16, 17, 20, 21, 22, 31, 35]. Existing learning-based methods generally fall into two categories: Graph comparison approaches compress node representations into holistic graph embeddings. In contrast, node comparison approaches estimate edit costs between node representations and iteratively refine mappings by down-weighting high-cost matches. Although graph comparison methods offer computational efficiency, they lose critical node-level details essential for modeling edit operations. In contrast, node comparison methods retain finer structural information, making them more suited to capture the flexibility of GED.
Figure 1: The black dashed lines are matches, and the red markings are edits. $f_{1}$ and $f_{2}$ are mappings. $f_{1}$ incurs 2 node substitutions (ns), 1 node insertion, 1 edge deletion, and 2 edge insertions (ei), denoted as $[2,1,0,1,2]$, while $f_{2}$ incurs $[4,1,0,0,1]$. When ns costs 2, other operations cost 1, $f_{1}$ is optimal with a cost of 8, while $f_{2}$ costs 10. However, if ei costs 2, others cost 1, $f_{1}$ costs 8 while $f_{2}$ costs 7, making $f_{2}$ the optimal choice.
Nevertheless, existing node comparison approaches exhibit two core limitations that directly violate the above principles of GED computation. First, they neglect the impact of operation costs during mapping determination. While GraphEdX [13] incorporates these costs when computing the final GED, it ignores them during mapping establishment, resulting in static mappings that do not adapt to cost variations. Second, these methods rely on node-level similarity scores to guide matching decisions, ignoring the graph-level dependencies between node matches, which reflect how matches jointly affect global structural alignment. Without modeling them, mappings are optimized in isolation, often resulting in myopic matches and necessitating costly post hoc refinement procedures, which substantially increase computational overhead.
In this paper, we propose Graph Edit Network (Gen), a novel framework that addresses these limitations by revisiting the fundamental principles of GED computation. Gen introduces two key innovations: (1) it incorporates operation costs prior to mapping establishment, ensuring that mapping decisions dynamically adapt to cost variations, and (2) it explicitly models graph-level dependencies between matches through a proactive guidance optimization strategy. By globally estimating and propagating node alignment difficulties across graphs, Gen captures the joint impact of individual node matches, facilitating informed one-shot matching and eliminating the need for costly refinements.
The contributions of this work are threefold: (i) We introduce a principled yet simple approach that integrates operation costs prior to mapping determination, addressing a fundamental gap in learning-based graph similarity models and improving adaptability to varying operation cost settings. (ii) We propose a proactive guidance optimization strategy, utilizing global alignment difficulty estimation and difficulty propagation to capture graph-level matching dependencies, enabling accurate one-shot matching without iterative refinements. (iii) Extensive empirical evaluations demonstrate the effectiveness, robustness of Gen, achieving substantial improvements in both accuracy and efficiency over existing methods.
2 Related Work
Graph similarity computation (GSC) aims to approximate measures such as graph edit distance (GED) [4, 16] using neural network-based approaches. This work focuses on end-to-end GED approximation methods, which can be further divided into graph-level and node-level comparison approaches. For more details, we refer the reader to recent surveys on this topic [18].
Graph-level comparison methods embed entire graphs into latent spaces and compute similarity based on graph representations [16, 21, 35, 22, 3]. GMN-emb [16] employs a gated mechanism to highlight important nodes during graph embedding. EGSC [21] introduces a teacher-student framework to balance prediction accuracy and storage efficiency. Eric [35] enforces node-graph alignment during training, and detaches it to accelerate inference. Greed [22] models the triangle inequality among graph embeddings by leveraging bounds on GED. To improve interpretability, TaGSim [3] predicts graph edit vectors (GEVs) representing edit operation counts. However, graph-level comparison methods lack sensitivity to mapping variations, which may result in limited accuracy under varying cost settings.
Node-level comparison methods compares node embeddings to capture fine-grained node similarities [16, 4, 5, 9, 20, 13]. SimGNN [4] augments graph representations with non-differentiable histogram features derived from node similarities, which cannot be directly learned during the training process. GMN-match [16] iteratively fuses cross-graph node representations using distance-based local attention, which may lead to suboptimal, myopic updates. GraphSim [5] reorders multi-scale similarity matrices via BFS traversal, which may oversimplify complex structures and reduce robustness. Recent approaches introduce combinatorial optimization to refine mappings. GOTSim [9] augments the similarity matrix to account for node deletion and insertion operations, then uses a linear assignment solver to align nodes. GEDGNN [20] and GraphEdX [13] leverage Gumbel-Sinkhorn networks for differentiable mapping, with GraphEdX further designing operation-specific distance measures. Nevertheless, these methods primarily operate in matrix form and often ignore graph-level matching dependencies, which may lead to suboptimal mappings that fail to reflect the true graph structural differences. Moreover, all existing methods compute static mappings independent of operation costs, limiting their adaptability to varying cost settings.
3 Background
[…]
Graph Edit Distance (GED). Given a source graph $S$ and a target graph $T$, the graph with fewer nodes was padded using dummy nodes to ensure $|\mathcal{V}_{S}|=|\mathcal{V}_{T}|$. GED is defined as the minimum total edit cost over all bijective mappings $f\in\mathcal{B}:\mathcal{V}_{S}\mapsto\mathcal{V}_{T}$, where the bipartite graph $\mathcal{B}=(\mathcal{V}_{S}\cup\mathcal{V}_{T},\mathcal{E}_{\mathcal{B}})$ represents the solution space of $f$, and each edge $m\in\mathcal{E}_{\mathcal{B}}$ denotes a potential node correspondence between $S$ and $T$. The GED from $S$ to $T$ is computed as:
[…]
Total $\text{\emph{Total Cost}: }c(S,T)_{f}=\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\text{c}(s,t)+\sum_{e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}}\text{c}(e,\epsilon)+\sum_{e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})}\text{c}(\epsilon,e^{\prime});$ (2)
[…]
In Eq. (1), $\text{c}(i_{s},i_{t})$ denotes the cost of editing element $i_{s}$ into $i_{t}$ via operation $op_{(i_{s},i_{t})}$, where $\text{cost}(\cdot)$ is a non-negative, customizable cost setting that specifies the operation cost, and $d(i_{s},i_{t})$ quantifies the attribute or structural dissimilarity between elements, typically 0 or 1 for individual nodes or edges. This formulation explicitly highlights the influence of operation costs in determining optimal mappings. In Eq. (2), $\mathcal{A}_{S}$ and $\mathcal{A}_{T}$ encode the intra-graph connections within the original graphs, and $\epsilon$ denotes a nonexistent edge. The terms $e\in f(\mathcal{A}_{S})\setminus\mathcal{A}_{T}$ and $e^{\prime}\in\mathcal{A}_{T}\setminus f(\mathcal{A}_{S})$ represent edges whose existence is not preserved under the mapping $f$, capturing the inter-graph discrepancies. Together, Eq. (2) models the graph-level dependencies between matches, reflecting how node correspondences jointly affect global structural alignment. Finally, in Eq. (3), $\text{GED}(S,T)$ is defined as the minimum of such total costs over all possible mappings.
4 Analyzing Flexible GED Computation in Embedding Space
Given graph pairs as input, learning-based node-level comparison methods approximate GED in the embedding space. They generally have three stages: (1) Encoding stage: they project the local structure of node onto the embedding space with an injective function $\Phi:\mathbf{X}\mapsto\mathbb{E}^{|\mathcal{V}|\times d}$, where $d$ is the dimensionality of the embeddings, $|\mathcal{V}|$ is the cardinality of $\mathcal{V}$; (2) Scoring stage: based on a distance/similarity measure function $\Psi$, they evaluate the distance for all possible source-target node embeddings; (3) Aligning stage: they use the pairwise embedding distances to guide the selection of matches $\mathbf{1}$, then use the total cost of a learned mapping to approximate GED.
In this work, we focus on the Scoring stage and the Aligning stage, as they are directly responsible for flexible and efficient GED computation. Existing methods typically apply cost-agnostic distance functions in the Scoring stage and use isolated node-level distances to guide the mapping establishment in the Aligning stage. However, these practices overlook the fundamental GED computation principles, namely operation cost sensitivity and graph-level dependencies between matches. Consequently, the resulting mappings are static and often suboptimal, necessitating costly post hoc refinements.
Limitation of Cost-agnostic Distance. Existing methods either ignore operation costs or consider them only at the end of the Aligning stage, treating them as irrelevant to the determination of optimal mappings. This results in static mappings that capture only intrinsic element dissimilarities. This limitation can be theoretically revealed by Eq. (1): When the cost setting function $\text{cost}(\cdot)$ assigns uniform operation cost (e.g., 1) to all operations, the total cost of a mapping corresponds to the number of edits. Let $f^{*}$ denote the optimal mapping under this uniform setting. When $\text{cost}(\cdot)$ varies, the total cost of $c(S,T)_{f^{*}}$ also changes, often rendering $f^{*}$ suboptimal, as shown in Figure 1. Consequently, existing methods cannot adapt to varying cost settings, limiting both the theoretical expressiveness and practical applicability of learned graph similarity models.
Limitation of Node-level Guidance. With the edit cost of pairwise node embeddings, the total cost of a learned mapping is given by:
$c(S,T)=\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\mathbf{1}_{s,t}\cdot c(s,t).$ (4)
Figure 2: Models marked with ‘*’ perform only a single round of refinement.
Where $\mathbf{1}_{s,t}$ denotes the (soft) selection of node correspondences. Existing methods typically determine $\mathbf{1}_{s,t}$ based on node-level distances. A representative example is GMN [16]. It derives local attention weights from pairwise node distances to guide inter-graph node representation fusions, and then refines node representations based on graph structures. Another representative approach is GraphEdX [13], which optimizes soft mappings via Gumbel-Sinkhorn networks, minimizing transport cost defined over node-level distances in a matrix form.
Both methods follow what we term reactive refinement strategies. They passively refine mappings based on local similarity signals without considering the graph-level dependencies between matches. This practice often leads to suboptimal intermediate mappings and costly refinements. The detailed comparison of methods can be found in Appendix A. Figure 2 provides empirical evidence that, when the number of refinement iterations in these models is limited to one, their MAE increases by 15% to 89% compared with the original settings recommended by the authors. This performance drop highlights their reliance on costly iterative refinements. We argue that efficient node-level GED approximation requires a proactive matching strategy that adheres to the principles of GED computation. Such a strategy should explicitly account for graph-level dependencies from the outset, capturing the joint impact of node matches on global structural alignment, thereby improving alignment quality and reducing computational overhead.
5 The Proposed GEN Architecture
Figure 3: Overview of Gen. The dashed node represents the dummy node.
An overview of the proposed Gen framework is illustrated in Figure 3. Given a source-target graph pair and customized costs per operation type for each possible match as inputs, Gen first uses a shared GNN Encoder as $\Phi$ to embed both graphs into a unified representation space, ensuring semantic alignment between node embeddings. Building on the analysis in Section 4, Gen introduces two components: the Cost-sensitive Scoring module $\Psi$ and the Proactive Guidance Optimization strategy, which jointly guide the soft selection of matches $\mathbf{1}$. $\Psi$ dynamically integrates operation costs with embedding distances, enabling adaptation to varying cost settings. The second module proactively optimizes guidance to capture the graph-level matching dependencies. It begins by estimating each node’s alignment difficulty to ensure global comparability. Then it propagates these difficulties within and across graphs via a Difficulty Propagation mechanism to capture intra-graph connections and inter-graph discrepancies. Gen uses the resulting guidance to construct $\mathbf{1}$, enabling informed one-shot matching and minimizing the need for reactive refinements. Finally, Gen computes the total cost by summing the weighted edit costs. The pseudo code of Gen is available in Appendix B.
5.1 Cost-sensitive Scoring
The node representation for node $v$ is initialized as the one-hot encoding of the node label, $x_{v}\in\mathbb{R}^{|\mathcal{X}|}$, which $|\mathcal{X}|$ is the number of label types. Gen first employs a $k$-layer GNN as $\Phi$ to generate node representations $\mathbf{X}^{k}$ based on the $k$-hop local structures. Due to the space limitation, the details of node feature initialization and the encoding process of $\Phi$ are provided in Appendix B.1. Given the initial label encoding $\mathbf{X}^{0}$ and node representations $\mathbf{X}^{k}$, this module computes the edit cost between all possible node pairs. For each node $v$, Gen first concatenate its $\mathbf{x}_{v}^{0}$ and $\mathbf{x}_{v}^{k}$: $\mathbf{h}_{v}=\text{Concat}(\mathbf{x}_{v}^{0},\mathbf{x}_{v}^{k})$, to emphasize label mismatches, where $\mathbf{h}_{v}\in\mathbb{R}^{|\mathcal{X}|+d}$.
Following Eq. (1), the edit cost from $s\in\mathcal{V}_{S}$ to $t\in\mathcal{V}_{T}$ is the operation cost-weighted inherent distance between $s$ and $t$ in the embedding space. We compute the inherent distance as $\text{d}(s,t)=\mathbf{h}_{s}-\mathbf{h}_{t}$. Instead of calculating distance involved in each operation type, Gen simplifies computation by using a unified $\text{d}(s,t)$ that represents all elements that should be edited. Given the predefined operation costs $cost{(s,t)}\in\mathbb{R}^{|op|}$ for each possible match, the edit cost is estimated as:
$c(s,t)=\text{Estimator}\big{(}\text{Concat}(cost{(s,t)},d(s,t))\big{)},$ (5)
where Estimator is a trainable multilayer perceptron (MLP). This design avoids naïvely applying each operation cost to the entire distance vector, which may contain elements unrelated to certain operations. By leveraging fully connected layers, the MLP can learn how different operation types interact with the latent differences in embeddings and produce a cost-aware correction.
Notably, Gen is the first method to incorporate operation costs before aligning, ensuring that while the inherent distances between embeddings remain consistent, they can be dynamically adjusted to accommodate varying cost settings. This is theoretically aligned with the GED computation principle in Eq. (1), enhancing the model’s understanding of GED computation. Further validation is provided in Experiment 2, highlighting the limitations of existing methods on this issue and the effectiveness of our solution.
5.2 Proactive Guidance Optimization
Unlike previous work that relies on the insufficient node-level distance as guidance, and reactively refines the match selections $\mathbf{1}$. Gen proactively optimize the guidance to capture the graph-level matching dependencies, constructing $\mathbf{1}$ in a single step.
Guidance Initialization. Edit costs between embeddings can vary significantly across matches. While the absolute cost of each match may be clear, comparing these costs directly can be misleading, as some matches might seem optimal for one node but not contribute optimally to the overall graph alignment, leading to myopic, locally optimal decisions for individual nodes. To address this, Gen proposes constructing global comparability. It first normalizes the edit costs into global cost ratios:
$\gamma(s,t)=\frac{\exp(c(s,t))}{\sum_{s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}}\exp(c(s,t))},$ (6)
where $\gamma(s,t)\in\mathbb{R}^{d}$ represents the cost ratio of a match compared with the overall cost of all candidate matches. Intuitively, a match with a higher matching cost can have a larger ratio, reflecting its global significance within the graph-level alignment. Each node is involved in different candidate matches, forming a spectrum of differences between the node and the other graph. Gen gathers the cost ratios for each involved node to estimate the node alignment difficulty.
$\tilde{\gamma}(s)$ $=\sum_{i\in\mathcal{V}_{T}}\gamma(s,i),\quad\tilde{\gamma}(t)=\sum_{i\in\mathcal{V}_{S}}\gamma(i,t).$ (7)
A higher alignment difficulty $\tilde{\gamma}(\cdot)$ indicates that accurate matching of the corresponding node is crucial to minimizing overall graph edit cost, providing a comprehensive assessment of each node’s relative importance.
Difficulty Propagation. Matching a node constrains the feasible matches of others, as neighborhood consistency must be preserved to minimize structural differences. Thus, the alignment difficulty of a node not only reflects its own ambiguity but also influences its neighbors. To capture such intra-graph connections, Gen propagates alignment difficulties within each graph:
$\hat{\gamma}(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Sum}\big{(}\tilde{\gamma}(v),\sum_{\mathcal{A}(v,u)=1}\tilde{\gamma}(u)\big{)}\bigg{)},$ (8)
where $\mathcal{N}(v)$ denotes the neighborhood of $v$. To strengthen the connection between difficulties and representations, Gen further refines $\hat{\gamma}(\mathcal{N}(v))$ using $\mathbf{h}_{v}$ as residual information:
$\gamma(\mathcal{N}(v))=\text{MLP}\bigg{(}\text{Concat}\big{(}\hat{\gamma}(\mathcal{N}(v)),\mathbf{h}_{v}\big{)}\bigg{)}.$ (9)
Here, $\gamma(\mathcal{N}(v))\in\mathbb{R}^{d}$ is the neighborhood alignment difficulty for each node $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T})$, broadening the decision-making scope from individual nodes to their local contexts.
A match connects two nodes within different graphs, each node can have a distinct alignment difficulty. Finally, for each node pair $(s,t)$, Gen considers such inter-graph discrepancies to balancing the alignment difficulties of the involved nodes. It computes certainty scores to approximate $\mathbf{1}_{s,t}$:
$\mathbf{1}_{s,t}=\text{MLP}\left(\text{Concat}(\gamma(s),\gamma(t))\right),$ (10)
where $\mathbf{1}_{s,t}\in\mathbb{R}^{d}$ reflects the soft confidence in matching $s$ and $t$, enabling more informed decisions in a single step by considering the graph-level dependencies between matches. Notably, all steps in this strategy are efficiently implemented using scatter operations over the sparse graph structure, which are well-optimized for GPU parallelism.
5.3 Prediction and Training Objective
Prediction. Given the $\mathbf{1}_{s,t}$ for each match, the edit cost per node is computed as:
$c(s)=\sum_{i\in\mathcal{V}_{T}}\mathbf{1}_{s,i}\cdot c(s,i),\quad c(t)=\sum_{i\in\mathcal{V}_{S}}\mathbf{1}_{i,t}\cdot c(i,t).$ (11)
Finally, the total cost of mappings is calculated as the sum of certainty-weighted edit costs, according to Eq. (4). The total costs for mappings established for graphs $S$ and $T$ are:
$c(S,T)_{f_{S}}=\sum_{s\in\mathcal{V}_{S}}c(s),\quad c(S,T)_{f_{T}}=\sum_{t\in\mathcal{V}_{T}}c(t).$ (12)
The predicted GED is obtained as:
$\text{GED}(S,T)=\text{Linear}\left(\text{Mean}\left(c(S,T)_{f_{S}},c(S,T)_{f_{T}}\right)\right),$ (13)
where $\text{Linear}(\cdot)$ reduces the dimensionality from $d$ to $1$. This process indirectly encourages consistent mappings between the $S$ and $T$, or the average matching cost increases accordingly.
Training Objective. Given the predicted GED $\hat{y}$ and the ground-truth GED $y$, Gen minimizes the Huber loss [12], which is computed as follows.
$\mathcal{L}(\hat{y},y)=\begin{cases}\frac{1}{2}(y-\hat{y})^{2}&\text{if }|y-\hat{y}|\leq\delta,\\ \delta\left(|y-\hat{y}|-\frac{1}{2}\delta\right)&\text{otherwise},\end{cases}$ (14)
where $\delta$ controls the transition between quadratic and linear behaviors, ensuring stable training under diverse error scales.
5.4 Complexity Analysis
The complexity of the GNN Encoder is $\mathcal{O}(k|\mathcal{E}|d_{\text{hidden}})$, where $k$ is the number of GNN layers and $\mathcal{E}$ is the set of edges. The complexity of Cost-sensitive Scoring is $\mathcal{O}(\max(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}\cdot d)$. The Proactive Optimization Strategy introduces an additional complexity of $\mathcal{O}(2(\max(|\mathcal{V}_{S}|,|\mathcal{V}_{T}|)^{2}d+|\mathcal{E}|d))$.
Table 1: Effectiveness of models on test sets. We bold the best and underline the second models. ‘OOM’ indicates out of memory.
Methods AIDS700nef LINUX IMDBMulti
RMSE $\downarrow$ MAE $\downarrow$ $\rho$uparrow $\tau$uparrow p@10 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $\rho$uparrow $\tau$uparrow p@10 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$ $\rho$uparrow $\tau$uparrow p@10 $\uparrow$
VJ $9.278_{\pm\text{0.000}}$ $8.709_{\pm\text{0.000}}$ $0.422_{\pm\text{0.000}}$ $0.296_{\pm\text{0.000}}$ $0.269_{\pm\text{0.000}}$ $3.407_{\pm\text{0.000}}$ $2.692_{\pm\text{0.000}}$ $0.728_{\pm\text{0.000}}$ $0.574_{\pm\text{0.000}}$ $0.934_{\pm\text{0.000}}$
Beam $5.111_{\pm\text{0.000}}$ $4.380_{\pm\text{0.000}}$ $0.486_{\pm\text{0.000}}$ $0.345_{\pm\text{0.000}}$ $0.371_{\pm\text{0.000}}$ $3.315_{\pm\text{0.000}}$ $2.841_{\pm\text{0.000}}$ $0.734_{\pm\text{0.000}}$ $0.575_{\pm\text{0.000}}$ $0.445_{\pm\text{0.000}}$
SimGNN $1.036_{\pm\text{0.040}}$ $1.075_{\pm\text{0.084}}$ $0.847_{\pm\text{0.008}}$ $0.670_{\pm\text{0.008}}$ $0.897_{\pm\text{0.005}}$ $0.621_{\pm\text{0.076}}$ $0.434_{\pm\text{0.066}}$ $0.919_{\pm\text{0.009}}$ $0.760_{\pm\text{0.012}}$ $0.869_{\pm\text{0.065 }}$ $13.721_{\pm\text{2.163}}$ $6.863_{\pm\text{1.161}}$ $0.899_{\pm\text{0.006}}$ $0.781_{\pm\text{0.006}}$ $0.978_{\pm\text{0.001}}$
GMN-match $1.068_{\pm\text{0.032}}$ $0.811_{\pm\text{0.021}}$ $0.873_{\pm\text{0.006}}$ $0.701_{\pm\text{0.008}}$ $0.899_{\pm\text{0.004}}$ $\underline{0.267_{\pm\text{0.007}}}$ $0.126_{\pm\text{0.008}}$ $\underline{0.946_{\pm\text{0.001}}}$ $0.805_{\pm\text{0.002}}$ $0.967_{\pm\text{0.009}}$ $\underline{8.401_{\pm\text{2.061}}}$ $\underline{3.694_{\pm\text{0.925}}}$ $\textbf{0.938}_{\pm\textbf{0.007}}$ $\textbf{0.843}_{\pm\textbf{0.017}}$ $\underline{0.985_{\pm\text{0.004}}}$
GMN-emb $1.645_{\pm\text{0.081}}$ $1.145_{\pm\text{0.033}}$ $0.854_{\pm\text{0.006}}$ $0.679_{\pm\text{0.006}}$ $0.793_{\pm\text{0.025}}$ $0.334_{\pm\text{0.033}}$ $0.174_{\pm\text{0.010}}$ $0.944_{\pm\text{0.001}}$ $0.803_{\pm\text{0.002}}$ $0.974_{\pm\text{0.014}}$ $12.340_{\pm\text{1.403}}$ $5.296_{\pm\text{0.707}}$ $0.926_{\pm\text{0.005}}$ $0.820_{\pm\text{0.010}}$ $0.966_{\pm\text{0.034}}$
GraphSim $2.017_{\pm\text{0.793}}$ $1.519_{\pm\text{0.580}}$ $0.382_{\pm\text{0.381}}$ $0.298_{\pm\text{0.306}}$ $0.365_{\pm\text{0.447}}$ $2.283_{\pm\text{1.008}}$ $1.721_{\pm\text{0.800}}$ $0.206_{\pm\text{0.371}}$ $0.175_{\pm\text{0.319}}$ $0.276_{\pm\text{0.359}}$ $159.571_{\pm\text{0.012}}$ $70.917_{\pm\text{0.002}}$ $0.003_{\pm\text{0.000}}$ $0.002_{\pm\text{0.000}}$ $0.001_{\pm\text{0.000}}$
GOTSim $1.584_{\pm\text{0.347}}$ $1.198_{\pm\text{0.249}}$ $0.691_{\pm\text{0.183}}$ $0.525_{\pm\text{0.149}}$ $0.783_{\pm\text{0.058}}$ $1.560_{\pm\text{0.621}}$ $1.242_{\pm\text{0.460}}$ $0.666_{\pm\text{0.363}}$ $0.528_{\pm\text{0.281}}$ $0.550_{\pm\text{0.066}}$ $129.623_{\pm\text{13.795}}$ $52.248_{\pm\text{7.199}}$ $0.611_{\pm\text{0.250}}$ $0.492_{\pm\text{0.208}}$ $0.474_{\pm\text{0.201}}$
TaGSim $1.113_{\pm\text{0.058}}$ $0.844_{\pm\text{0.039}}$ $0.841_{\pm\text{0.012}}$ $0.663_{\pm\text{0.014}}$ $0.896_{\pm\text{0.009}}$ $0.678_{\pm\text{0.066}}$ $0.438_{\pm\text{0.051}}$ $0.913_{\pm\text{0.010}}$ $0.758_{\pm\text{0.012}}$ $0.909_{\pm\text{0.029}}$ $11.565_{\pm\text{3.310}}$ $5.451_{\pm\text{1.340}}$ $0.873_{\pm\text{0.053}}$ $0.764_{\pm\text{0.059}}$ $0.979_{\pm\text{0.008}}$
ERIC $1.259_{\pm\text{0.316}}$ $0.971_{\pm\text{0.228}}$ $0.782_{\pm\text{0.097}}$ $0.605_{\pm\text{0.095}}$ $0.836_{\pm\text{0.111}}$ $0.271_{\pm\text{0.130}}$ $\underline{0.120_{\pm\text{0.104}}}$ $0.944_{\pm\text{0.007}}$ $0.803_{\pm\text{0.012}}$ $0.973_{\pm\text{0.027}}$ $9.866_{\pm\text{0.662}}$ $4.462_{\pm\text{0.292}}$ $0.932_{\pm\text{0.005}}$ $0.834_{\pm\text{0.008}}$ $0.977_{\pm\text{0.006}}$
Greed $0.900_{\pm\text{0.016}}$ $0.705_{\pm\text{0.012}}$ $0.885_{\pm\text{0.001}}$ $0.715_{\pm\text{0.002}}$ $0.930_{\pm\text{0.005}}$ $0.441_{\pm\text{0.003}}$ $0.327_{\pm\text{0.002}}$ $0.939_{\pm\text{0.001}}$ $0.791_{\pm\text{0.001}}$ $0.960_{\pm\text{0.002}}$ $19.974_{\pm\text{20.988}}$ $6.948_{\pm\text{5.880}}$ $0.890_{\pm\text{0.039}}$ $0.778_{\pm\text{0.047}}$ $0.837_{\pm\text{0.303}}$
GEDGNN $1.607_{\pm\text{0.618}}$ $1.148_{\pm\text{0.354}}$ $0.747_{\pm\text{0.114}}$ $0.577_{\pm\text{0.104}}$ $0.739_{\pm\text{0.155}}$ $0.313_{\pm\text{0.098}}$ $0.188_{\pm\text{0.071}}$ $0.945_{\pm\text{0.005}}$ $\underline{0.806_{\pm\text{0.007}}}$ $\underline{0.993_{\pm\text{0.005}}}$ $15.743_{\pm\text{9.533}}$ $7.070_{\pm\text{3.471}}$ $0.897_{\pm\text{0.036}}$ $0.785_{\pm\text{0.046}}$ $0.934_{\pm\text{0.076}}$
GraphEdX $\textbf{0.735}_{\pm\textbf{0.015}}$ $\underline{0.581_{\pm\text{0.011}}}$ $\underline{0.910_{\pm\text{0.003}}}$ $\underline{0.749_{\pm\text{0.004}}}$ $\underline{0.951_{\pm\text{0.003}}}$ $0.378_{\pm\text{0.135}}$ $0.222_{\pm\text{0.081}}$ $0.941_{\pm\text{0.006}}$ $0.797_{\pm\text{0.011}}$ $0.973_{\pm\text{0.021}}$ OOM OOM OOM OOM OOM
Ours (Gen) $\underline{0.738_{\pm\text{0.007}}}$ $\textbf{0.577}_{\pm\textbf{0.005}}$ $\textbf{0.915}_{\pm\textbf{0.002}}$ $\textbf{0.756}_{\pm\textbf{0.003}}$ $\textbf{0.963}_{\pm\textbf{0.003}}$ $\textbf{0.166}_{\pm\textbf{0.009}}$ $\textbf{0.084}_{\pm\textbf{0.008}}$ $\textbf{0.948}_{\pm\textbf{0.000}}$ $\textbf{0.810}_{\pm\textbf{0.001}}$ $\textbf{0.995}_{\pm\textbf{0.004}}$ $\textbf{7.531}_{\pm\textbf{0.293}}$ $\textbf{3.492}_{\pm\textbf{0.128}}$ $\underline{0.937_{\pm\text{0.002}}}$ $\underline{0.839_{\pm\text{0.002}}}$ $\textbf{0.987}_{\pm\textbf{0.001}}$
6 Evaluations
In this section, we evaluate Gen against twelve baselines, including two conventional inexact algorithms, on three commonly used GED datasets and synthetic graphs, focusing on Effectiveness, Time Efficiency, Adaptation to Varying Cost Settings, and Performance on Larger Synthetic Graphs. The hyperparameter senstitivity analysis are available in Appendix D.4.
6.1 Experimental Setup
Dataset. We conduct experiments on three commonly used GED datasets: AIDS700nef, Linux and IMDBMulti. Following prior work [16, 3, 20], we generate synthetic graphs with a node size ranging from 100 to 400, as extreme data to to assess model performance on larger graphs. We generate 51,200 graphs for training and another 38,400 graphs in three node size groups: $[100,200]$, $[200,300]$, and $[300,400]$, for testing. The detailed dataset descriptions can be found in Appendix C.1.
Baselines. We include representative, recent, and state-of-the-art GED computation models and approximation algorithms as baselines: (1) Graph Comparison Methods: GMN-emb [16], TaGSim [3], Greed [22] and Eric [35]; (2) Node Comparison Method: GMN-match [16], SimGNN [4], GraphSim [5], GOTSim [9], GEDGNN [20] and GraphEdX [13]. (3) Approximate Algorithms: Volgenant–Jonker (VJ) algorithm [10] and Beam [19]. We generally use the official implementations and hyperparameters provided by the authors. More details about the baselines and the implementation of Gen can be found in Appendix C.2 and C.3.
Experimental Protocol. We partitioned the datasets into training and test graphs at a 4:1 ratio, with 20% of the training graphs used for validation. Training pairs are sampled from the training set. All models are trained for ten warm-up epochs and evaluated on the validation set every epoch. Early stopping with a patience of 50 is applied to prevent overfitting and ensure that each model is sufficiently trained. More details are available in Appendix C.4.
Evaluation Metrics. We evaluate models’ prediction error and ranking ability. For error metrics, we report Root Mean Square Error (RMSE) and Mean Absolute Error (MAE). For ranking, we use Spearman’s Rank Correlation Coefficient ($\rho$), Kendall’s Rank Correlation Coefficient ($\tau$), and Precision at $k$ ($p@k$). Details of metric definitions are provided in Appendix C.5. We report the average and standard deviation of results across five runs.
6.2 Results
Effectiveness. We compare the performance of Gen against all baselines under a uniform cost setting (operation costs set to 1). Results for VJ and Beam on IMDBMulti are excluded, as the training targets of this dataset are the smallest distances produced by these algorithms. As shown in Table 1, Gen consistently ranks first or second across all metrics and datasets, demonstrating strong overall performance. GNN-based methods generally outperform conventional ones due to their ability to encode richer structural information. Notably, graph comparison methods (e.g., Greed, Eric) outperform most node comparison approaches, suggesting no definitive advantage between the two paradigms under this setting. Gen clearly outperforms most node comparison methods, as they often rely on isolated node-level distances, while Gen proactively optimizes guidance to capture graph-level matching dependencies, enabling informed one-shot matching. Consequently, Gen achieves up to a 37.8% reduction in RMSE and a 30% reduction in MAE on LINUX, and improves ranking performance by up to 4.1%, with only marginal declines ($\leq$0.4%) in a few cases. Further discussions and and a ranking case study can be found in Appendix D.1 and D.2, respectively.
Table 2: ‘Rank’ indicates the current ranking based on RMSE and show the improvements over the ranking for the AIDS700nef dataset in Table 1 under a uniform cost setting, as shown in parentheses.
Dataset Setting 1 Setting 2
[…]
SimGNN $1.923_{\pm\text{0.082}}$ $1.487_{\pm\text{0.064}}$ $0.859_{\pm\text{0.010}}$ $0.680_{\pm\text{0.013}}$ 5 (-1) $1.100_{\pm\text{0.024}}$ $0.862_{\pm\text{0.015}}$ $0.955_{\pm\text{0.002}}$ $0.820_{\pm\text{0.003}}$ 4 (0)
GMN-match $4.874_{\pm\text{0.054}}$ $3.421_{\pm\text{0.054}}$ $0.423_{\pm\text{0.017}}$ $0.295_{\pm\text{0.014}}$ 11 (-6) $4.761_{\pm\text{0.152}}$ $3.317_{\pm\text{0.086}}$ $0.471_{\pm\text{0.009}}$ $0.344_{\pm\text{0.009}}$ 11 (-6)
GMN-emb $5.742_{\pm\text{0.309}}$ $4.070_{\pm\text{0.110}}$ $0.448_{\pm\text{0.009}}$ $0.311_{\pm\text{0.006}}$ 13 (-3) $8.694_{\pm\text{0.117}}$ $6.093_{\pm\text{0.097}}$ $0.182_{\pm\text{0.028}}$ $0.125_{\pm\text{0.020}}$ 13 (-3)
GraphSim $3.874_{\pm\text{1.636}}$ $2.707_{\pm\text{1.047}}$ $0.356_{\pm\text{0.419}}$ $0.283_{\pm\text{0.336}}$ 9 (+2) $3.226_{\pm\text{1.574}}$ $2.488_{\pm\text{1.212}}$ $0.380_{\pm\text{0.458}}$ $0.320_{\pm\text{0.388}}$ 9 (+2)
GOTSim $3.316_{\pm\text{0.211}}$ $2.635_{\pm\text{0.186}}$ $0.768_{\pm\text{0.058}}$ $0.577_{\pm\text{0.053}}$ 8 (0) $2.754_{\pm\text{0.049}}$ $2.173_{\pm\text{0.034}}$ $0.760_{\pm\text{0.021}}$ $0.564_{\pm\text{0.017}}$ 8 (0)
TaGSim $2.120_{\pm\text{0.051}}$ $1.605_{\pm\text{0.017}}$ $0.844_{\pm\text{0.003}}$ $0.662_{\pm\text{0.003}}$ 6 (0) $1.188_{\pm\text{0.081}}$ $0.864_{\pm\text{0.032}}$ $0.950{\pm\text{0.004}}$ $0.816_{\pm\text{0.005}}$ 6 (0)
ERIC $2.319_{\pm\text{0.173}}$ $1.826_{\pm\text{0.116}}$ $0.784_{\pm\text{0.035}}$ $0.599_{\pm\text{0.033}}$ 7 (0) $1.587_{\pm\text{0.138}}$ $1.250_{\pm\text{0.096}}$ $0.911_{\pm\text{0.011}}$ $0.747_{\pm\text{0.014}}$ 7 (0)
Greed $4.607_{\pm\text{0.045}}$ $3.396_{\pm\text{0.040}}$ $0.494_{\pm\text{0.012}}$ $0.348_{\pm\text{0.009}}$ 10 (-7) $5.874_{\pm\text{0.072}}$ $4.200_{\pm\text{0.060}}$ $0.237_{\pm\text{0.013}}$ $0.170_{\pm\text{0.010}}$ 12 (-9)
GEDGNN $1.782_{\pm\text{0.921}}$ $1.331_{\pm\text{0.598}}$ $0.916_{\pm\text{0.082}}$ $0.770_{\pm\text{0.104}}$ 4 (+5) $1.102_{\pm\text{0.630}}$ $0.850_{\pm\text{0.466}}$ $0.953_{\pm\text{0.049}}$ $0.830_{\pm\text{0.080}}$ 5 (+4)
GraphEdX $1.249_{\pm\text{0.014}}$ $0.978_{\pm\text{0.008}}$ $0.934_{\pm\text{0.002}}$ $0.786_{\pm\text{0.002}}$ 3 (-2) $0.714_{\pm\text{0.012}}$ $0.567_{\pm\text{0.009}}$ $0.977_{\pm\text{0.001}}$ $0.871_{\pm\text{0.002}}$ 3 (-2)
Ours (Gen) $\underline{1.061_{\pm\text{0.032}}}$ $\underline{0.771_{\pm\text{0.033}}}$ $\underline{0.952_{\pm\text{0.003}}}$ $\underline{0.823_{\pm\text{0.006}}}$ 2 (0) $\underline{0.545_{\pm\text{0.007}}}$ $\underline{0.409_{\pm\text{0.007}}}$ $\underline{0.984_{\pm\text{0.000}}}$ $\underline{0.893_{\pm\text{0.001}}}$ 2 (0)
TaGSim mix $7.106_{\pm\text{0.137}}$ $5.249_{\pm\text{0.159}}$ $0.307_{\pm\text{0.024}}$ $0.217_{\pm\text{0.017}}$ 14 (new) $4.088_{\pm\text{0.171}}$ $3.087_{\pm\text{0.182}}$ $0.510_{\pm\text{0.014}}$ $0.382_{\pm\text{0.011}}$ 10 (new)
GraphEdX mix $5.024_{\pm\text{0.411}}$ $4.081_{\pm\text{0.357}}$ $0.620_{\pm\text{0.058}}$ $0.446_{\pm\text{0.047}}$ 12 (new) $12.918_{\pm\text{1.144}}$ $10.919_{\pm\text{1.334}}$ $0.218_{\pm\textbf{0.051}}$ $0.142_{\pm\textbf{0.038}}$ 14 (new)
Ours (Gen mix) $\textbf{0.898}_{\pm\textbf{0.044}}$ $\textbf{0.642}_{\pm\textbf{0.043}}$ $\textbf{0.979}_{\pm\textbf{0.002}}$ $\textbf{0.881}_{\pm\textbf{0.006}}$ 1 (new) $\textbf{0.479}_{\pm\textbf{0.013}}$ $\textbf{0.353}_{\pm\textbf{0.015}}$ $\textbf{0.988}_{\pm\textbf{0.001}}$ $\textbf{0.905}_{\pm\textbf{0.003}}$ 1 (new)
Time Efficiency. We evaluate the training and inference time of all models. Results for VJ and Beam are excluded, as prior studies have shown that learning-based methods significantly outperform them in efficiency [4, 31]. As shown in Figure 5, our proposed Gen ranks fourth or fifth fastest in training time, outperforming all node comparison (NC) models and closely approaching the slowest graph comparison (GC) baseline, Eric. While SimGNN shows similar batch runtimes, it requires more training epochs to converge. At inference, GC methods are generally more time-efficient than NC ones, the latter exhibiting at least quadratic complexity. Compared with the most relevant and competitive NC baseline, GMN-match, Gen reduces inference time by up to 72.7% while achieving comparable or better performance. This efficiency gain stems from our proactive optimization strategy, which captures graph-level dependencies and reduces the need for iterative refinements.
Adaptation to Varying Costs. We assess the adaptability of models to non-uniform and varying cost settings using the node-labeled AIDS700nef dataset. We evaluate two cost settings: Setting 1 assigns costs of 1 for node substitution, 2 for node deletion, 1 for node insertion, 3 for edge deletion, and 1 for edge insertion, denoted as [1, 2, 1, 3, 1]. In Setting 2, the corresponding costs are [3, 2, 3, 0, 2]. As shown in Table 2, (i) under non-uniform cost settings: Gen achieves the best performance in both settings, showing strong adaptability to asymmetric edit costs. In contrast, Greed performs poorly due to its use of symmetric distance measure, and GMN-match suffers from error accumulation caused by its greedy local fusion mechanism. Notably, rank improvements are more prominent among NC models compared with GC ones, indicating NC models’ suitability for flexible GED estimation. (ii) Under varying costs: We further train Gen mix, TaGSim mix, and GraphEdX mix on a dataset combining both settings. Gen mix substantially outperforms all baselines, demonstrating better adaptability. The sharp performance degradation of TaGSim mix and GraphEdX mix highlights their key limitation, as they treat operation costs as irrelevant during aligning stage, they produce static mappings that are suboptimal under non-uniform settings. In contrast, the improvement of Gen mix over Gen highlights the robustness and generalization capability of our principled solution. This indicates that exposure to diverse cost settings does not impair Gen’s performance, but instead facilitates a deeper understanding of GED computation, leading to more accurate predictions.
Table 3: Ablations Study on our proposed components.
Dataset AIDS700nef LINUX IMDBMulti
RMSE $\downarrow$ p@10 $\uparrow$ RMSE $\downarrow$ p@10 $\uparrow$ RMSE $\downarrow$ p@10 $\uparrow$
Gen wo global $0.808_{\pm\text{0.038}}$ $0.954_{\pm\text{0.009}}$ $0.178_{\pm\text{0.013}}$ $0.993_{\pm\text{0.005}}$ $9.415_{\pm\text{0.977}}$ $0.972_{\pm\text{0.023}}$
Gen wo dependencies $1.101_{\pm\text{0.041}}$ $0.905_{\pm\text{0.006}}$ $0.296_{\pm\text{0.034}}$ $\underline{0.990_{\pm\text{0.005}}}$ $14.391_{\pm\text{2.568}}$ $0.982_{\pm\text{0.004}}$
Gen wo intra-graph $\underline{0.746_{\pm\text{0.011}}}$ $\underline{0.958_{\pm\text{0.002}}}$ $\textbf{0.159}_{\pm\textbf{0.015}}$ $\textbf{0.995}_{\pm\textbf{0.003}}$ $\underline{8.439_{\pm\text{0.656}}}$ $\underline{0.984_{\pm\text{0.005}}}$
Gen $\textbf{0.738}_{\pm\textbf{0.007}}$ $\textbf{0.963}_{\pm\textbf{0.003}}$ $\underline{0.166_{\pm\text{0.009}}}$ $\textbf{0.995}_{\pm\textbf{0.004}}$ $\textbf{7.531}_{\pm\textbf{0.293}}$ $\textbf{0.987}_{\pm\textbf{0.001}}$
Ablation Study. We evaluate the effectiveness of our proposed strategy. We test three variants: (1) Gen wo global: This variant replaces the global cost ratio with absolute node matching costs to estimate alignment difficulties. (2) Gen wo dependencies: This variant reweighs matches solely based on their global cost ratios, where higher ratios result in lower weights. (3) Gen wo intra-graph: This variant does not consider intra-graph connections. Note that the inter-graph discrepancies cannot be ablated separately as they produce the matching decisions. As shown in Table 3, Gen wo dependencies greedily reweighs matches based on global cost ratios, leading to a significant performance drop across all datasets, with RMSE increasing by up to 91.1%. On the other hand, Gen wo global incurs a smaller increase, at most 25%, indicating that considering dependencies is more crucial than the global cost ratio. Meanwhile, ignoring the difficulties of neighboring nodes has a smaller impact. Gen wo intra-graph showcased comparable performance on AIDS700nef and LINUX, yet incurred a 12.1% increase in RMSE on IMDBMulti. This is likely because the matches in IMDBMulti are more ambiguous, making the neighborhood alignment difficulty a crucial factor.
(a) Training Time.
(b) # Epochs.
(c) Inference Time.
Figure 4: Time Efficiency of GSC models. The bars with black edges represent node comparison-based methods.
Figure 5: Performance of models on larger synthetic graphs across diverse node size groups.
Performance on Larger Synthetic Graphs. GED is typically applied to small graphs, as its exact computation becomes challenging for graphs with more than 16 nodes [6]. To evaluate the performance of GSC models under extreme conditions, we followed [20] and assessed the performance of GSC models on larger synthetic graphs. We compared GED predictions among baseline methods and our proposed Gen against ground-truth GEDs. Figure 5 illustrates the results for the best NC and GC models during this experiment, GEDGNN and TaGSim, against our proposed Gen. The results show that Gen achieved the most accurate predictions, with data points closely clustered around the line representing perfect alignment between predictions and GEDs across all node size groups. The full results can be found in Appendix D.3.
7 Conclusion
This paper presents Graph Edit Networks (Gen), a learning-based model that captures the inherent flexibility of graph edit distance (GED) computation with an efficient proactive optimization strategy. We highlight a key limitation in existing methods in handling flexible GED computation: they neglect or treat operation costs as irrelevant when determining the mappings, resulting in static mappings across cost settings. To address this issue, we follow the GED computation principle and propose incorporating operation costs before establishing mappings. Instead of reactively refining mappings with insufficient node-level distances, we introduce a novel proactive optimization strategy. This strategy enables effective guidance capturing graph-level matching dependencies, allowing accurate one-shot match selections. Experiments on both real-world and synthetic datasets validate the model’s effectiveness, efficiency, and adaptability. Our proposed approach can be applied in graph retrieval scenarios, where Gen can return the top $k$ relevant graphs based on a query or rerank candidate graphs generated by fast graph comparison models to improve precision.
Appendix A Comparison Table
To further highlight the key differences in guidance and refinement strategies, we summarize the comparisons in Table 4.
Appendix B Pseudo Code for Gen
The pseudo code for Gen can be found in Algorithm 1.
Algorithm 1 The Gen Architecture
0:  A padded graph pair $(S,T)$ each has N nodes, where $S=(\mathbf{X}_{S},\mathcal{A}_{S})$, query graph $T=(\mathbf{X}_{T},\mathcal{A}_{T})$, the customized operation cost matrix $\mathcal{C}\in\mathcal{R}^{N^{2}\times|op|}$
0:  The estimated GED value between $S$ and $T$
1:  $\mathbf{X}^{0}_{S}=\operatorname{Linear}_{pre}(\mathbf{X}_{S})\in\mathbb{R}^{|\mathcal{V}|_{S}\times d_{hidden}}$ Initialize Node Embedding
2:  $\mathbf{X}^{0}_{T}=\operatorname{Linear}_{pre}(\mathbf{X}_{T})\in\mathbb{R}^{|\mathcal{V}|_{T}\times d_{hidden}}$
3:  for $l=1,\dots,K$ do
4:     $\mathbf{X}_{S}^{l}=\operatorname{GNN}(\mathbf{x}_{s}^{l-1}\in\mathbf{X}_{S}^{l-1},\mathcal{A}_{S})\in\mathbb{R}^{|\mathcal{V}_{S}|\times d_{hidden}}$ GNN Encoder
5:     $\mathbf{X}_{T}^{l}=\operatorname{GNN}(\mathbf{x}_{t}^{l-1}\in\mathbf{X}_{T}^{l-1},\mathcal{A}_{T})\in\mathbb{R}^{|\mathcal{V}_{T}|\times d_{hidden}}$
6:  end for
7:  $\mathbf{X}_{S}=\operatorname{MLP}(\mathbf{X}_{S}^{k})\in\mathbb{R}^{|\mathcal{V}_{S}|\times d}$ Reduce Dimensionality
8:  $\mathbf{X}_{T}=\operatorname{MLP}(\mathbf{X}_{T}^{k})\in\mathbb{R}^{|\mathcal{V}_{T}|\times d}$
9:  $\mathbf{H}_{S}=\text{Concat}(\mathbf{X}_{S},\mathbf{X}_{S}^{k})\in\mathbb{R}^{|\mathcal{V}_{S}|\times(|\mathcal{X}|+d)}$
10:  $\mathbf{H}_{T}=\text{Concat}(\mathbf{X}_{T},\mathbf{X}_{T}^{k})\in\mathbb{R}^{|\mathcal{V}_{T}|\times(|\mathcal{X}|+d)}$
11:  $\text{dist}(S,T)=\mathbf{H}_{S}-\mathbf{H}_{T}\in\mathbb{R}^{N^{2}\times(|\mathcal{X}|+d)}$ Cost-aware Scoring
12:  $c(S,T)=\text{Estimator}(\text{Concat}(\mathcal{C},\text{dist}(S,T)))\in\mathbb{R}^{N^{2}\times d}$
13:  for $c(s,t)\in c(S,T)$ do
14:     $\gamma(s,t)=exp(c(s,t))/c(S,T)\in\mathbb{R}^{d}$ Global Cost Ratios Initialization
15:  end for
16:  for $s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}$ do
17:     $\tilde{\gamma}(s)=\sum_{i\in\mathcal{V}_{T}}\gamma(s,i)\in\mathbb{R}^{d}$ Node Alignment Difficulty Computation
18:     $\tilde{\gamma}(t)=\sum_{i\in\mathcal{V}_{S}}\gamma(i,t)\in\mathbb{R}^{d}$
19:  end for
20:  for $v\in(\mathcal{V}_{S}\cup\mathcal{V}_{T}$do
21:     $\hat{\gamma}(v)=\text{MLP}\bigg{(}\text{Sum}(\tilde{\gamma}(v),\text{Sum}(\tilde{\gamma}(v^{\prime}):\mathcal{A}(v,v^{\prime})=1))\bigg{)}\in\mathbb{R}^{d}$ Neighborhood Alignment Difficulty Computation
22:  end for
23:  $\gamma(S)=\text{MLP}\bigg{(}\text{Concat}(\hat{\gamma}(S),\mathbf{H}_{S})\bigg{)}\in\mathbb{R}^{|\mathcal{V}_{S}|\times d}$ Add Residual Info
24:  $\gamma(T)=\text{MLP}\bigg{(}\text{Concat}(\hat{\gamma}(T),\mathbf{H}_{T})\bigg{)}\in\mathbb{R}^{|\mathcal{V}_{T}|\times d}$
25:  $\xi(S,T)=\text{MLP}\bigg{(}\text{Concat}(\gamma(S),\gamma(T))\bigg{)}\in\mathbb{R}^{N^{2}\times d}$ Compute Certainty
26:  for $s\in\mathcal{V}_{S},t\in\mathcal{V}_{T}$ do
27:     $c(s)=\sum_{i\in\mathcal{V}_{T}}\xi(s,i)\cdot c(s,i)\in\mathbb{R}^{d}$ Compute Certainty-weighted Cost
28:     $c(t)=\sum_{i\in\mathcal{V}_{s}}\xi(i,t)\cdot c(i,t)\in\mathbb{R}^{d}$
29:  end for
30:  $c(S)=\sum_{s\in\mathcal{V}_{S}}c(s)\in\mathbb{R}^{d}$ Compute The Total Matching Costs
31:  $c(T)=\sum_{t\in\mathcal{V}_{T}}c(t)\in\mathbb{R}^{d}$
32:  $\text{GED}(S,T)=\text{Linear}\bigg{(}\text{Mean}(c(S),c(T))\bigg{)}\in\mathbb{R}^{1}$
33:  return $\text{GED}(S,T)$ Predict GED
B.1 GNN Encoder
Given a graph as input, Gen uses a $k$-layer GNNs to generate representations for each node based on its $k$-hop local structure. The node representation for node $v$ is initialized as the one-hot encoding of the node label, $\mathbf{x}_{v}\in\mathbb{R}^{|\mathcal{X}|}$, which $|\mathcal{X}|$ is the number of label types. Nodes in the unlabeled graphs can be regarded as having an identical label. We pad a smaller graph between $S$ and $T$ with dummy nodes to ensure $|\mathcal{V}_{S}|=|\mathcal{V}_{T}|$. The representations of dummy nodes are initialized as zeros. We first smooth $\mathbf{x}_{v}$ with a Linear layer to obtain $\mathbf{x}_{v}^{0}\in\mathbb{R}^{d_{hidden}}$. Based on the adjacency matrix $\mathcal{A}$ of graphs, GNNs update node information for each node with aggregated information from the neighborhoods. The encoding process at layer $l$ for node $v$ can be represented as follows.
$\mathbf{x}^{l}_{v}=\psi\bigg{(}\mathbf{x}^{l-1}_{v},\phi(\mathbf{x}^{l-1}_{u}:\mathcal{A}(v,u)=1)\bigg{)}$ (15)
Where $\phi$ is an aggregation function, and $\psi$ is a combination function, typically permutation-invariant operations like Sum, Mean, or Max. $\mathbf{x}^{l}_{v}$ and $\mathbf{x}^{l-1}_{v}\in\mathbb{R}^{d_{hidden}}$ represent the output of node $v$ from layers $l$ and $l-1$, respectively. $\mathbf{x}^{l-1}_{u}\in\mathbb{R}^{d_{hidden}}$ represents the output of $v$’s neighbor at the previous layer. The final layer representation $\mathbf{x}^{k}_{v}$ is then passed through a Multilayer Perceptron (MLP) to reduces the dimensionality of the representation to $d$. By altering GNNs with the ones considering edge information, this method can be extended to edge-labeled graphs without losing generalization.
Appendix C Experimental Setup
C.1 Dataset
We conduct experiments over three commonly used GED datasets:
  • AIDS700nef: 700 node labeled chemical compound structures with ten or less than ten nodes;
  • Linux: 1,000 unlabeled Program Dependence Graphs (PDG) generated from the Linux kernel, with ten or less than ten nodes;
  • IMDBMulti: 1500 unlabeled ego-networks of movie actors/actresses, with up to 89 nodes.
The graphs within the datasets and the exact or approximate (for IMDBMulti dataset) GED values between all possible graph pairs under a uniform cost setting are collected and provided by [4], resulting in abundant graph pairs. With the same graphs in the AIDS700nef dataset, we compute the GEDs under two non-uniform settings with the A*-based algorithm [2]. Note that exact GED computation is already challenging for graphs with 16 nodes [6]. Additionally, we generate synthetic graphs, among Erdős-Rényi, power-law, and Watts–Strogatz small-world graphs with a node size ranging from 100 to 400, as extreme data to further evaluate the effectiveness of models under the impact of graph size. For each synthetic graph, we apply $\Delta$ distinct operations to generate a target graph. The number of operations ranges from 10-30 percent of edge numbers. Following previous work [16, 3, 20], we regard this $\Delta$ as the ground-truth GED for synthetic graph pairs. We generate 51,200 graphs as training sets and another 38,400 graphs within three node size groups, including $[100,200]$, $[200,300]$, and $[300,400]$, as test sets. The detailed statistics of graphs and distributions of GEDs of these datasets can be found in Table 5 and Figure 6, respectively.
Figure 6: The GED distributions of datasets.
Table 5: Statistics of datasets.
Dataset #Graphs #Pairs #Testing Pairs #Labels Avg. #nodes Avg. #edges Max. #nodes Max. #edges Min. #nodes Min. #edges
AIDS700nef 700 490K 78,4K 29 8.90 8.80 10.00 14.00 2.00 1.00
LINUX 1000 1M 160k 1 7.58 6.94 10.00 13.00 4.00 3.00
IMDBMulti 1500 2.25M 360k 1 13.00 65.94 89.00 1467.00 7.00 12.00
Synthetic 51,200 25,6 K - 1 230.24 1789.02 400.00 8552.00 100.00 100.00
[100, 200] 12,800 6.4K - 1 148.61 1328.51 200.00 5806.00 100.00 196.00
[200, 300] 12,800 6.4K - 1 247.45 1979.85 300.00 7638.00 200.00 396.00
[300, 400] 12,800 6.4K - 1 348.14 2295.15 400.00 8083.00 300.00 100.00
C.2 Baselines
We include representative, recent, and state-of-the-art GED computation models as baselines: (1) Graph Comparison Methods: GMN-emb [16], TaGSim [3], Greed [22] and Eric [35]; (2) Node Comparison Method: GMN-match [16], SimGNN [4], GraphSim [5], GOTSim [9], GEDGNN [20], GraphEdX [13]. Additionally, we compare with well-known approximate algorithms, such as the Volgenant–Jonker (VJ) algorithm [10] and Beam [19]. Hungarian [23] is not included as it has been incorporated in GOTSim. Specifically:
  • VJ [10]: Volgenant–Jonker (VJ) algorithm is a linear sum assignment that greedily performs bipartite graph matching to minimize costs between nodes of two graphs.
  • Beam [19]: Beam search employs a best-first strategy, expanding partial solutions with the lowest-cost matches at each step to approximate GED.
  • SimGNN [4]: a representative node comparison-based method that generates histogram features based on pairwise node similarity scores.
  • GMN-match [16]: a representative node comparison-based method that alternatively propagates node information within and between graphs to update node information and implicitly establish node mapping.
  • GMN-emb [16]: a representative graph comparison-based method which generated graph representation using gate-weighted node representations.
  • GraphSim [5]: a representative node comparison-based method that reorders the similarity matrix by performing BFS traversal to capture the local affinity of matches.
  • GOTSim [9]: a representative node comparison-based method that leverages linear assignment solver to explicitly establish node mapping at each layer.
  • TaGSim [3]: a representative operation type-aware graph-level comparison method.
  • ERIC [35]: the state-of-the-art graph comparison-based method proposed to train models with node-graph similarity in an unsupervised manner, and decouple this module in the inference.
  • Greed [22]: a method handles GED computation by computing the Euclidean distance between graph representations.
  • GEDGNN [20]: the recent node comparison-based method that explicitly builds node mapping by optimizing the graph transport costs.
  • GraphEdX [13]: the state-of-the-art node comparison-based method that defines a family of distance measures for each operation types and reweighs the total cost with operation costs.
We generally use the official implementation and hyperparameters provided by the authors. The implementations of VJ and Beam are sourced from [29]. For models predicting the normalized GED value, we remove their last sigmoid layer to ensure a fair comparison.
C.3 Implementation Details
We apply a 5-layer Graph Isomorphism Network (GIN) [28] with jump connections as GNN encoder, whose hidden state and output are of 64 dimensionality. All MLPs in this work have adopted layer normalization and applied ReLU as the activation function. The outputs of these MLPs are of a dimensionality of 32. We used Adam as the optimizer, with a fixed learning rate of 0.001.
C.4 Experimental Protocol
All models are trained on a NVIDIA A6000 GPU with an AMD EPYC 7282 CPU. The batch size is set to 128 graph pairs for training and 2048 for validation and test. Due to the large number of possible graph pairs, each epoch consists of 100 iterations, except for synthetic graphs where pair counts are limited. Each model undergoes ten warm-up epochs and is evaluated on the validation set at the end of each epoch. Early stopping with a patience of 50 is applied, terminating training if the validation loss does not improve over 50 consecutive validation steps, to prevent overfitting while ensuring sufficient training.
C.5 Evaluation Metrics
Following previous work, we evaluate models based on prediction error and ranking ability. Prediction error measures the point-wise difference between predicted and ground-truth GED values. At the same time, ranking ability is assessed by simulating retrieval scenarios in graph databases, where training graphs serve as data graphs, test graphs serve as queries. During the testing phase, each model calculates the GED between the query graph and every graph in the database. In this context, the model should return the most relevant data graphs with the lowest GED for each query. We use Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) to assess GED prediction errors. Similar to MSE, RMSE is sensitive to outliers but presents errors in the original units for better interpretability, while MAE captures general errors.
For ranking performance, we use Spearman’s Rank Correlation Coefficient ($\rho$), Kendall’s Rank Correlation Coefficient ($\tau$), and Precision at $k$ ($p@k$). $\rho$ is computed as:
$\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)},$
where $d_{i}$ represents the difference in ranks for each data point between the two rankings, and $n$ is the total number of data points, the factor of 6 is included to normalize the sum of squared rank differences, ensuring that $\rho$ falls within the range $[-1,1]$ [25]. This metric measures the strength and direction of monotonic relationships. $\tau$ evaluates ordinal consistency and is computed as:
$\tau=\frac{C-D}{\binom{n}{2}},$
where $C$ is the number of pairs ranked in the same order in both rankings, $D$ is the number of pairs ranked in opposite orders in the two rankings, $\binom{n}{2}$ is the total number of possible pairs in a set of $n$ items. $p@k$ checks the relevance of the top $k$ predicted graphs against the ground truth GED ranking for each query. It calculates the intersection of the predicted top $k$ data graphs and the ground-truth top $k$ results, dividing the count of the intersection by $k$ for each query, and then averaging the results across all queries. To evaluate the ranking ability, we sorted all data graphs for each query based on both the ground truth and predicted GED values in ascending order, and then compared the resulting rankings.
Appendix D Results
D.1 Effectiveness
We provide additional analysis of the baseline models’ performance across datasets:
  • GraphEdX performs well on the node-labeled AIDS700nef dataset by leveraging XOR operations as node-level similarity. However, it underperforms on the unlabeled LINUX dataset and runs out of memory on IMDBMulti due to its design choice of padding all graphs to the same size, rather than within graph pairs like GEDGNN and Gen.
  • GMN-match fuses and propagates node representations across graphs in an iterative manner. While it performs competitively on AIDS700nef and LINUX, it suffers from error accumulation caused by its greedy fusion process.
  • GraphSim employs BFS-based matrix reordering to improve structural alignment. This design, however, introduces instability, as reflected by large standard deviations across runs.
Figure 7: Case Study of ranking results on AIDS700nef (labeled) and IMDBMulti (unlabeled).
D.2 Case Study
Figure 7 presents a case study comparing the top-ranked graphs retrieved by Gen and the ground truth GED rankings computed by A*. The example demonstrates that Gen closely approximates the true ranking, successfully identifying and ranking structurally similar graphs.
D.3 Performance on Larger Synthetic Graphs
Additional experimental in Figure 8 show that GMN-match produced higher outlier predictions. While SimGNN’s predictions generally align with the true distribution of GED, they are less concentrated than ours and exhibit more outlier predictions. We did not report results for another graph comparison method, Eric, because it proved particularly challenging to train due to gradient exploding issues. Despite our efforts to implement layer normalization at each layer of Eric, these attempts did not alleviate the problem. Aligned with other graph comparison models, Greed does not perform as well as its node comparison counterparts.
Figure 8: Performance of models on larger synthetic graphs across diverse node size groups.
D.4 Hyperparameter Sensitivity
We analyze Gen’s sensitivity to key hyperparameters such as GNN backbones, number of layers, and dimensionality. As shown in Figure 9, GIN proves to be the optimal backbone for Gen, consistently achieving the lowest RMSE. Adding layers generally improves performance, with the effect varying across datasets. On AIDS700nef and LINUX, performance increases steadily with more layers, while on IMDBMulti, it peaks at four layers for RMSE reduction. Similarly, increasing dimensionality typically reduces RMSE, but on IMDBMulti, performance worsens beyond 32/16 dimensions, likely due to overfitting.
(a) AIDS700nef
(b) LINUX
(c) IMDBMulti
Figure 9: Hyperparameter Sensitivity